### 05 Mar 2026 v1

# Efficient Reasoning

Abstract
Recent advances in large language models (LLMs) have enabled agentic systems for sequential decision-making. Such agents must perceive their environment, reason across multiple time steps, and take actions that optimize long-term objectives. However, existing web agents struggle on complex, long-horizon tasks due to limited in-context memory for tracking history, weak planning abilities, and greedy behaviors that lead to premature termination. To address these challenges, we propose StructuredAgent, a hierarchical planning framework with two core components: (1) an online hierarchical planner that uses dynamic And/Or trees for efficient search and (2) a structured memory module that tracks and maintains candidate solutions to improve constraint satisfaction in information-seeking tasks. The framework also produces interpretable hierarchical plans, enabling easier debugging and facilitating human intervention when needed. Our results on WebVoyager, WebArena, and custom shopping benchmarks show that StructuredAgent improves performance on long-horizon web-browsing tasks compared to standard LLM-based agents.
1 Introduction
Large language models (LLMs) have recently achieved remarkable performance across a wide range of natural language processing tasks, giving rise to agentic LLMs: language models augmented with the ability to perceive, reason, and act within interactive environments. These agentic LLMs are increasingly being deployed in diverse real-world applications such as enterprise workflow automation, customer support, software development, and autonomous research assistance.
A particularly promising application area is the development of web-browsing agents, which navigate and interact with the internet to accomplish complex, goal-driven tasks including information seeking, form filling, and transactional activities. Despite notable progress, building robust and interpretable web-based agents remains challenging.
Although many web tasks are intuitively straightforward for humans, even state-of-the-art web agents powered by advanced LLMs continue to struggle with complex, multi-step web-browsing tasks [Yang et al., 2025, Erdogan et al., 2025, Qi et al., 2025]. These limitations can be largely traced to four main challenges: limited in-context memory, which restricts the agent’s ability to track alternative solution paths across multiple web pages, especially as dense web pages can exceed 20k tokens (see Figure˜2) and most existing agents consider only the current page and action history; an inability to construct and execute complex plans over extended horizons; a lack of robust error-handling mechanisms to recover from failures and explore alternative solutions [Koh et al., 2025]; and greedy decision-making that can lead to premature task termination without satisfying all user constraints [Schmied et al., 2025] (see Appendix˜E).
In cognitive science [Newell et al., 1972], human problem solving is framed as heuristic search coupled with means-ends analysis, where complex tasks are recursively decomposed into manageable subproblems. This perspective motivates hierarchical planning in AI, where agents improve efficiency by breaking down complex goals into nested subgoals and actions at multiple levels of abstraction. Such methods are particularly well-suited to open-ended, web-based tasks characterized by uncertainty and expansive state and action spaces. However, designing effective admissible heuristics for these domains remains an open challenge.
Recent research has explored explicit planning mechanisms for web agents. Koh et al. [2025] use tree-based planning with A* search to explore alternative solution paths upon failure; however, their method operates at the level of individual actions, which limits exploitation of compositional task structure, and the absence of admissible heuristics can lead to inefficient path switching. Yang et al. [2025] propose incremental tree construction with prune-and-branch strategies, but rely solely on the LLM for plan construction and execution, often resulting in simplistic plans and limited error recovery. Erdogan et al. [2025] introduce dynamic plan revisions, but permit only sequential plans determined prior to execution, without error back-propagation.
Research Gap. Despite these advances, current web-based agents still lack the adaptability, robustness, and interpretability that characterize human problem solving.
To address these challenges, we propose StructuredAgent, a hierarchical planning framework that interleaves planning and execution, enabling agents to dynamically construct and execute ordered And/Or hierarchical planning trees during task execution. Our framework decomposes complex tasks into hierarchies of nodes (subgoals), Or nodes (alternative strategies), and atomic web actions, empowering agents to reason about fallback options and generate interpretable plans that transparently reflect their internal decision-making. In contrast to conventional LLM agents that require the model to generate and manage entire plans, our approach separates responsibilities: the framework constructs and maintains the planning tree, while the LLM is invoked only for local tree operations such as node expansion or repair. Our framework uses a modified greedy depth-first search for tree expansion and supports dynamic plan revision and error back-propagation, enabling agents to efficiently revise or prune parts of a plan as new information becomes available and to effectively explore large, complex search spaces. Furthermore, for information-seeking tasks in web environments, we introduce a structured memory module that tracks candidate entities and the constraints they satisfy, enhancing constraint satisfaction throughout planning and execution. Our main contributions are:
  • We introduce a hierarchical planning framework enabling inference-time planning via ordered And/Or trees, supporting compositional reasoning, dynamic plan revision, and error backpropagation for adaptive decision-making, while producing interpretable plans that facilitate error localization and human intervention.
  • For web-based information-seeking tasks, we develop a structured memory module that tracks candidate entities and their satisfied constraints, improving constraint satisfaction during planning and execution.
  • We demonstrate the effectiveness of our approach on the challenging WebVoyager and WebArena benchmarks, as well as a custom-designed shopping benchmark.
Figure 1: Illustration of StructuredAgent solving a web task via greedy DFS of a dynamically constructed And/Or tree. The root node represents the task objective and is expanded into subtasks that are progressively refined and executed. Node types are color-coded to distinguish OR ($\vee$), AND ($\wedge$), and ACTION nodes. The top half tracks the corresponding DFS stack states, where each node enters the stack in one of three states: ENTERING (ENT), EXITING (EXT), or FAILED (FAIL). Node repair and pruning are supported but not triggered in this example, and human intervention is optional.
2 Preliminaries
Web-browsing tasks are partially observable and inherently stochastic, as web page content and structure can change dynamically over time, either unpredictably or in response to the agent’s actions, and the agent lacks access to the full underlying browser state. The agent is typically limited to the current HTML page, prior actions, and extracted metadata, requiring it to make decisions under uncertainty while maintaining internal beliefs about the environment.
We model the agent-browser interaction as a Partially Observable Markov Decision Process (POMDP) $\langle\mathcal{O},\mathcal{S},\mathcal{A},R,T,p_{0},\gamma\rangle$ Kaelbling et al. [1998], where $o\in\mathcal{O}$ is an observation comprising the task description, agent-constructed plans, the HTML DOM, and interaction metadata; $s\in\mathcal{S}$ is the true underlying state of the browser and agent; and $a\in\mathcal{A}$ is an action, either an atomic browser operation (click, scroll, type, go back) or a token sequence describing a subgoal or strategy. The reward function $R\mathrel{\mathop{\ordinarycolon}}\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ assigns a scalar reward to each state-action pair, the transition function $T\mathrel{\mathop{\ordinarycolon}}\mathcal{S}\times\mathcal{A}\to\Delta^{|\mathcal{S}|}$ defines a distribution over next states, $p_{0}\in\Delta^{|\mathcal{S}|}$ is the initial state distribution, and $\gamma\in(0,1)$ is the discount factor. The agent’s policy $\pi\mathrel{\mathop{\ordinarycolon}}\mathcal{S}\to\Delta^{|\mathcal{A}|}$ maps states to distributions over actions and is, in general, history-dependent [Kaelbling et al., 1998]. We represent the agent’s policy using an LLM. At each timestep $t$, the observation history $h_{t}=\{o_{1},\dots,o_{t}\}$ is encoded as a token sequence $x_{t}=[x_{1},\dots,x_{m}]$ serving as the input prompt. The action $a_{t}=[y_{1},\dots,y_{n}]$ is then generated autoregressively as $\Pr(a_{t}\mid h_{t})=\prod_{i=1}^{n}\pi(y_{i}\mid x_{1},\dots,x_{m},y_{1},\dots,y_{i-1}).$
The StructuredAgent framework consists of two core components: (1) a StructuredMemory module that efficiently tracks intermediate decisions and candidate solutions during exploration, and (2) an And/Or Tree Planner that hierarchically decomposes tasks while reasoning over alternative strategies. Together, these components enable robust error isolation, dynamic replanning, and more reliable execution in complex web environments.
At the core of our approach is an explicit And/Or planning tree, which the agent constructs and maintains dynamically throughout the task. As new information is acquired, the agent continuously expands, revises, and prunes the planning tree to eliminate unpromising paths, enhancing computational efficiency while preventing unnecessary tree growth. Unlike conventional LLM agents, which lack the capability to autonomously construct and execute complex hierarchical plans, StructuredAgent separates planning responsibilities to improve both reliability and interpretability. Rather than delegating full tree construction and maintenance to the LLM, as in AgentOccam [Yang et al., 2025], StructuredAgent manages all aspects of tree construction directly, invoking the LLM only for targeted, well-defined operations including node expansion, repair, completion checking, and pruning. Constraining LLM queries to well-defined subtasks renders agent reasoning substantially more interpretable and robust. We provide detailed descriptions of these operations in Section˜3.3.
Next, we formalize the structure of the And/Or planning tree and describe its construction and traversal. We then detail the four tree operations and the observation and nodes aggregation module used to monitor task progress and agent state. Finally, we present the planning algorithm that integrates these components for efficient and adaptive execution.
(a) WebVoyager
(b) WebArena
Figure 2: Distribution of average HTML observation lengths (in tokens) per timestep across agent trajectories for WebVoyager and WebArena. Token counts reflect only raw web page content, excluding any additional context provided to the agent.
3 StructuredAgent
3.1 AND/OR Planning Tree Structure
The StructuredAgent constructs And/Or trees composed of three primary node types. (1) The node nodes represent conjunctive subgoals. All child nodes must succeed for the parent node to succeed. Depending on the task, the agent may or may not need to execute the children in a specific order. (2) Or nodes represent alternative strategies for achieving a subgoal. The parent Or succeeds as soon as any child node succeeds. (3) Action nodes serve as leaf nodes in the And/Or tree. Each Action node corresponds to a concrete browser-level operation such as typing a query, scrolling, clicking a button, navigating back or home, or recording a note.
This hierarchical structure allows the agent to decompose tasks into subgoals and reason explicitly about alternative solution paths. For example, when solving a task such as ‘‘find a coffee filter under $\mathdollar 50$‘‘, the agent might construct an node with two children representing searching for “coffee filter” and applying a price filter. Each subgoal can further decompose or branch into Or nodes that encode different navigation or filtering strategies. Both and Or nodes can expand into any combination of , Or, or Action nodes. This flexibility allows the agent to represent a wide variety of decompositions while capturing the sequential dependencies that arise in web-based workflows.
In Figure˜13 we highlight a benefit of our framework’s interpretable hierarchical planning structure, namely that it naturally supports human-in-the-loop oversight, whereby an operator can inspect the And/Or tree, detect incorrect subgoal decompositions, and inject corrective subgoals at any layer of the plan before execution proceeds further.
3.2 And/Or Tree Construction and Traversal
We now outline the main idea and key challenges in enabling an agent to perform dynamic planning using And/Or trees. Given a web-based task, the goal is to interleave planning and execution so that the agent can continually adapt its strategy as new information becomes available. To achieve this, we use an LLM as a high-level controller that issues localized tree-manipulation directives. These directives determine the type of a node, expand the node, repair it by adding or removing children, or verify whether its objective has been satisfied. The surrounding framework manages tree traversal and executes the corresponding operations.
We use a greedy and iterative depth-first search (DFS) to expand nodes in StructuredAgent, starting at the root node that represents the task objective. When the agent encounters an node, it processes its children sequentially according to the node’s prescribed ordering. When it encounters an Or node, the agent selects a single child to explore by querying an LLM for an estimate of its likelihood of success based on summaries of prior actions and observations. If the agent reaches an Action node, it executes the action in the environment and updates the tree with newly observed information. This update may require pruning or revising certain nodes.
Although DFS is traditionally applied to a fixed and finite graph, in our setting it serves as a mechanism for incrementally constructing and maintaining an evolving And/Or tree whose structure changes as the agent gathers new information. This distinction is essential in web-based environments, where the search space and the environment dynamics are large, stochastic, and difficult to model. As a consequence, subgoals generated during planning may be imprecise or infeasible. These imperfections often lead to node failures during execution. To address such failures, StructuredAgent must support systematic node revision and, after repeated unsuccessful attempts, node pruning. Pruning a child of an node further invalidates all subsequent siblings in its prescribed execution order, and the resulting structural changes must propagate upward through the tree to preserve global plan coherence. Maintaining this coherence may require revisiting and reprocessing affected nodes multiple times during traversal.
These requirements motivate two key modifications to standard DFS, where we first introduce three execution states for all nodes in the DFS stack: Entering, Exiting, and Failed. These states support upward propagation of structural changes and execution failures, allowing the framework to process a node multiple times as the tree evolves. A node enters the Entering state when it has not yet been processed or when some of its children remain unprocessed. It transitions to Exiting once its execution is complete for Action nodes, once all of its children have been processed for nodes, or once the selected child has been processed for Or nodes. In this state, the framework evaluates whether the node’s objective has been satisfied. A node enters the Failed state if execution fails for Action nodes or if any child fails or is pruned for /Or nodes. This state indicates that revision or pruning is required. Note that a node may appear in any of these states multiple times.
The second modification addresses the fact that classical iterative DFS uses only a binary notion of visited and unvisited nodes, which is insufficient in our setting because dynamic pruning and revision may require processing a node multiple times. To handle this, we assign each node one of six status values, summarized in Table˜6 (in the appendix). These values capture the node’s structural, execution, and failure conditions, enabling the framework to represent and manage the additional complexity introduced by dynamic tree evolution.
Figure 3: Illustration of the StructuredAgent Framework.
Figure 4: Illustration of node state transitions during iterative modified greedy depth-first search on AND-OR tree. Unlike traditional depth-first search, nodes may be revisited and repeatedly transition through Entering, Exiting, and Failed states.
We present the high-level tree construction and execution procedure used by StructuredAgent in Algorithm˜1. Full details of the planning algorithm are provided in Section˜D.1. We now describe the fundamental tree operations and the agent modules required by the StructuredAgent framework.
Figure 5: Overview of the primary operations used to construct and maintain the And/Or tree structure.
3.3 Tree Operations and Agent Modules
We now describe the tree structure used by our agent and the set of operations that construct and maintain it under the guidance of an LLM acting as a high-level controller. Recall that the web–task-solving problem is formulated as a Partially Observable Markov Decision Process (POMDP), requiring the agent’s policy to condition on the full history of observations. Because storing the entire observation history in the LLM’s in-context memory is infeasible, prior work advocates constructing concise summaries of browser interactions and using these summaries as proxies for the full historical state. Following this approach, the agent maintains several internal representations, detailed in Table˜5.
Building on these representations, we introduce a suite of specialized operators that utilize the associated metadata to enable dynamic decision-making and structured planning. The operators Node Expansion, Node Repair, Global Tree Update, Node Completion Check, and Observation Summarizer each prompt the backbone LLM to perform context-specific inference. The framework then verifies and executes the model’s predictions. See Section˜E.1 for detailed prompts, inputs, and outputs for all operators.
The Node Expansion operator determines node types and generates appropriate children: subgoals for nodes, executable actions for Action nodes, and candidate strategies with scores for Or nodes. The Node Repair operator revises failed And/Or nodes by proposing structural modifications or pruning nodes when recovery is no longer viable. Through Global Tree Update, the system maintains tractability by updating the global search tree, pruning irrelevant branches, and refining node descriptions as new context becomes available. The Node Completion Check operator evaluates whether nodes have met their objectives by analyzing child-node outcomes and relevant task summaries. Finally, the Observation Summarizer module produces internal representations of task progress and observations, enabling context-aware reasoning and effective planning. Further implementation details and operator specifications appear in Appendix˜B.
stack $\leftarrow[(root,\mathrm{ENTERING})]$
counter $\leftarrow 0$
while stack $\neq\emptyset$ and counter << BUDGET do
  (node,state) $\leftarrow$ stack.pop()
 if $(node.status=\mathrm{SUCCESS}\wedge node.revision\_count\geq\mathrm{MAX})\;\vee\;(node.status=\mathrm{DELETED})$ then
     continue
  else if $node.status=\mathrm{PRUNED}$ then
     BACKTRACK_FAILURE(node)
     // Alg. 5
     continue
 if $state=\mathrm{ENTERING}$ then
     stack,node $\leftarrow$ PROCESS_NODE_ENTERING(node,stack)
     // Alg. 2
  else if $state=\mathrm{EXITING}$ then
     stack,node $\leftarrow$ PROCESS_NODE_EXITING(node,stack)
     // Alg. D.1
  else if $state=\mathrm{FAILED}$ then
     stack,node $\leftarrow$ PROCESS_NODE_FAILED(node,stack)
     // Alg. 4
  end if
end while
Algorithm 1 AND_OR_TREE_CONSTRUCTION(root)
Equipped with the And/Or tree and core operators, we now describe the modified greedy, iterative DFS procedure that StructuredAgent uses to build, update, and execute its plan. The agent maintains a DFS stack and processes each node in one of Entering, Exiting, or Failed states.
When the agent encounters a node in the Entering state, it first checks whether the node has been expanded. If the node is unexpanded, the agent invokes Node Expansion to determine the node type and generate its children. These children consist of an ordered list of sub-goals for nodes, ranked alternative strategies for Or nodes, or a browser-level action for Action nodes. After expansion, the agent pushes the node back to the stack in the Exiting state and then pushes children in the Entering state. While all children are added in order for nodes, only the highest-ranked unprocessed child is considered for Or nodes. For an Action node, the agent immediately executes the corresponding browser operation, updates its internal context through the Observation Summarizer, and applies Global Tree Update to prune irrelevant branches and refine node descriptions.
When the agent processes a node in the Exiting state, it evaluates whether the node has satisfied its objective. An node is marked successful only if all valid children succeed. The Node Completion Check operator confirms that the children collectively satisfy the node’s description. An Or node is marked successful if any child has succeeded. An Action node is marked successful or failed depending on the outcome of the executed browser interaction. Once a node succeeds, the agent does not revisit it.
When the agent encounters a node in the Failed state, it begins recovery. If the failed node is an Action node, the agent prunes it. If it is part of an node, the remaining unexecuted siblings are removed because the conjunctive requirement can no longer be met. If an node fails, the agent checks its remaining revision budget. When possible, it calls Node Repair to add/modify children. If repair is unsuccessful or the budget is exhausted, the node and all of its descendants are pruned and the failure propagates upward. If an Or node fails, the agent first checks for any remaining unprocessed children. If such children exist, it selects the next most promising one. Otherwise, it attempts repair within the revision limit or prunes the node if repair is not viable. Pruning an Or node also propagates failure upward.
Through iterative node expansion, action execution, outcome evaluation, and targeted repair/pruning, this DFS procedure enables StructuredAgent to construct and adapt its hierarchical plan while incorporating new observations and responding to failures. See Appendix˜D for full algorithmic details.
Example. We illustrate the modified greedy DFS on the task ‘‘Find a vegan chocolate brownie recipe with 4+ rating on allrecipes.com’’ (see Figure˜3). The agent begins with only the root node $1$ on the stack in Entering. In this example, Observation Summarizer and Global Tree Update are invoked after each Action node execution ($1.1$, $1.2.1$, $1.3$); Observation Summarizer updates the internal context while Global Tree Update makes no updates to the tree.
Steps 1–2 [Expand & Execute root and $1.1$] Node Expansion identifies node $1$ as an node and generates three ordered children: $1.1$ (type query), $1.2$ (select recipe), and $1.3$ (note recipe), all of which must succeed. Node $1$ is re-pushed in Exiting and its children are pushed in Entering in prescribed order. Node $1.1$ is then popped and identified as an Action node; the search query executes immediately and node $1.1$ is marked success and transitions through Exiting.
Steps 3–4 [Expand & Execute $1.2$ and $1.2.1$] Node $1.2$ is popped in Entering and identified by Node Expansion as an Or node, encoding alternative recipe-selection strategies. Its children are $1.2.1$ (Select Recipe A, score $1.0$) and $1.2.2$ (Select Recipe B, score $0.95$). Since this is an Or node, only the highest-ranked child $1.2.1$ is pushed in Entering. Node $1.2.1$ is then popped and executed as an Action node (clicking Recipe A). Node Completion Check confirms the Or objective of $1.2$ is satisfied, so $1.2.2$ is never explored and node $1.2$ is marked success.
Steps 5–6 [Execute $1.3$] Node $1.3$ is popped in Entering and executed as an Action node (noting the recipe). Node $1.3$ succeeds and transitions through Exiting. The root node $1$ is then popped in Exiting, and Node Completion Check confirms that all three children have succeeded. The stack empties and the agent terminates with success.
No node enters Failed in this trace, so repair and pruning are not triggered. Had node $1.2$ exhausted both Or alternatives without success, it would enter Failed, causing its unexecuted sibling $1.3$ to be pruned and failure to propagate upward to root node $1$, which would then invoke Node Repair as long as its revision budget is not exhausted.
3.4 Structured Memory
Unstructured note-taking, as used in AgentOccam, fails to reliably retain items discovered during exploration. Specifically, agents may discard or overlook items encountered during earlier browsing steps, causing them to revisit already-rejected pages or fail to pursue alternative candidates when an initial choice turns out not to satisfy all constraints. We address this with a Structured Memory module designed for information-seeking tasks such as recommendation or research queries, where the agent must identify items matching a set of user-specified constraints. We refer to such items as candidate entities. The module extracts these constraints from the task description and organizes candidates in a dynamic table (See Figure˜12), where each row represents a candidate entity and columns capture constraints, attributes, or task-specific notes. The schema remains flexible, allowing new columns to be added as additional constraints or features are encountered during exploration. As the agent browses, it prompts the language model to add, update, or delete candidate entries based on the latest information. During decision-making in the Node Expansion and Node Repair modules, the top-$K$ candidates satisfying the most constraints are retrieved to guide subsequent actions, reducing redundant exploration throughout the task.
4 Experiments
Amazon Easy Amazon Hard WebVoyager Easy
Method kimi Gemini GPT Avg kimi Gemini GPT Avg Human kimi Gemini GPT Avg
k2 ($\uparrow$) 2.5F ($\uparrow$) 4.1 ($\uparrow$) k2 ($\uparrow$) 2.5F ($\uparrow$) 4.1 ($\uparrow$) k2 ($\uparrow$) 2.5F ($\uparrow$) 4.1 ($\uparrow$)
StructuredAgent 87.5 75.0 87.5 83.3 45.0 28.3 33.3 35.6 58.3 80.9 73.0 80.9 78.3
StructuredAgentMem 80.0 72.5 82.5 78.3 43.3 28.3 41.7 37.8 63.3 79.8 66.3 75.3 73.8
AgentOccam 75.0 62.5 70.0 69.2 36.7 16.7 31.7 28.4 56.3 84.3 71.9 83.2 79.8
BasicClaudeAction 77.5 60.0 77.5 71.7 28.3 18.3 23.3 23.3 0.4 74.2 67.4 70.8 70.8
SteP 62.5 57.5 62.5 60.8 13.3 10.0 13.3 12.2 - 74.2 60.7 69.7 68.2
Table 1: Success rates (%) across Amazon Easy, Amazon Hard, and WebVoyager Easy task sets. All methods use Claude 3.5 Sonnet as the backbone and are evaluated with three LLM judges: Kimi-k2, Gemini-2.5-Flash, and GPT-4.1. Human evaluation scores are additionally reported for Amazon Hard. Bold indicates the best performance per column. StructuredAgent achieves the highest average on Amazon Easy (83.3%), while StructuredAgentMem leads on Amazon Hard (37.8%).
Method Map Shopping Reddit GitLab Overall
($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$)
StructuredAgent 0.450 0.528 0.726 0.458 0.526
AgentOccam 0.463 0.380 0.601 0.467 0.464
BasicClaudeAction 0.370 0.303 0.303 0.281 0.310
WebarenaReplication 0.315 0.313 0.197 0.288 0.285
AgentOccam (GPT 4-o) 0.446 0.352 0.623 0.329 0.413
SteP (GPT 4-o) 0.333 0.269 0.564 0.225 0.321
Table 2: Success rates on WebArena across four task categories: Map (128), Shopping (192), Reddit (114), and GitLab (196). All methods use Claude 3.5 Sonnet unless otherwise noted (GPT suffix indicates GPT-4 backbone). Bold indicates the best performance per column. StructuredAgent achieves the highest overall score (52.6%), with particular strength on Shopping and Reddit tasks.
(a) GITLAB
(b) MAP
(c) REDDIT
(d) SHOPPING
Figure 6: Trajectory length distributions (number of browser actions) across three WebArena task categories. Each violin plot shows the distribution of trajectory lengths per method, with the mean marked.
(a) Amazon Easy
(b) Amazon Hard
Figure 7: Trajectory length distributions (number of browser actions) for Amazon Easy and Hard tasks, with the mean marked per violin. On harder tasks, StructuredAgent and StructuredAgentMem take substantially more actions than AgentOccam and BasicClaudeAction, reflecting the additional exploration required for complex multi-constraint shopping tasks.
Figure 8: Pairwise disagreement rates (%) between LLM judges and human evaluators across three task categories. Disagreement rate measures the percentage of tasks where two evaluators produce different outcomes (success vs. failure). Amazon Hard exhibits the highest disagreement, particularly between LLM judges and human evaluators, suggesting it is the most challenging to evaluate consistently.
Benchmarks. We conduct experiments on: (1) Custom Complex Shopping, a set of 60 long-horizon Amazon shopping tasks where the agent must identify products satisfying detailed natural-language constraints (typically requiring 10–30 interaction steps); (2) WebVoyager [He et al., 2024], a curated subset of 129 real-world tasks spanning diverse live websites; and (3) WebArena [Zhou et al., 2023], where we evaluate on 630 out of 812 tasks, excluding those associated with the ShoppingAdmin site and tasks where agents consistently fail due to website errors or evaluator failures. All main experiments use Claude 3.5 v2 and Claude 3.7 as backbone LLMs for StructuredAgent and all baselines unless otherwise specified. We provide full benchmark and baseline details in Section˜F.1.
Metrics. We evaluate all agents using task completion rate. For WebVoyager, we adopt an LLM-as-a-judge protocol using GPT-4.1, Kimi-k2-0905-preview, and Gemini-2.5-Flash as judges, with additional human evaluation on Amazon shopping tasks. For WebArena, we use the benchmark’s official code-based evaluators, supplemented by human annotation for cases where HTML- or URL-based evaluators may produce false negatives.
Baselines. We compare StructuredAgent against: (1) AgentOccam [Yang et al., 2025], (2) BasicClaudeAction, a Claude-based agent with full action history context, (3) SteP [Sodhi et al., 2024], a stack-based agent with human-written strategies, and (4) WebarenaReplication, the reference WebArena agent with CoT reasoning.
4.1 Results
RQ1) How does StructuredAgent compare with other LLM-based agents on short-horizon and complex long-horizon tasks? Table˜1 reports the success rates of all agents on Amazon Easy, Amazon Hard, and WebVoyager Easy tasks using Claude 3.5 as the backbone model. An LLM-as-a-Judge model considers a task successful if the agent satisfies all item-level and task-level constraints. StructuredAgent achieves the highest average performance on both Amazon Easy and Amazon Hard tasks, outperforming AgentOccam by 14% on Amazon Easy and by at least 7% on Amazon Hard tasks according to human evaluation. On WebVoyager Easy, StructuredAgent performs comparably to the baselines with a marginal drop of 1.5%, which we attribute to the relative simplicity of these tasks since they do not require the complex multi-step planning from which our framework most benefits.
We observe higher success rates with human evaluation than with LLM-as-a-Judge scores on Amazon Hard tasks. We attribute this discrepancy to the strict nature of automated evaluation, as Amazon Hard tasks require the agent to retrieve information about multiple items while satisfying several simultaneous constraints, and LLM-as-a-Judge models penalize partial satisfaction more harshly than human annotators do. Nevertheless, both evaluation protocols agree on the relative ordering of agents, confirming the robustness of the observed trends.
We further observe that augmenting StructuredAgent with the Structured Memory module (StructuredAgentMem) improves performance on Amazon Hard tasks by 5% under human evaluation, but slightly degrades performance on Amazon Easy and WebVoyager Easy. This behavior is expected because on simpler tasks that do not require tracking multiple candidate solutions, the Structured Memory module generates overly strict intermediate constraints that act as noise and degrade task completion. The Structured Memory module therefore works best for complex information-seeking tasks that require the agent to simultaneously track multiple items and satisfy diverse user constraints.
Fig. 8 reports the disagreement rate between pairs of LLM-as-a-Judge models and between human and automated evaluation. We observe substantially lower disagreement rates on easy tasks and notably higher disagreement rates on Amazon Hard tasks. Tasks that ask the agent to retrieve one or more bundles of items drive a primary source of this disagreement. Human evaluators typically accept a single valid bundle as a correct response, whereas LLM-as-a-Judge models expect multiple bundles due to perceived ambiguity in the task specification.
We also evaluate agents across major WebArena task categories, namely Map, Shopping, Reddit, and GitLab in Table˜2. Since these tasks require an average of 12 to 15 steps and are not primarily information-seeking in nature, we compare all baselines against StructuredAgent without the Structured Memory module. StructuredAgent outperforms all other baselines on Shopping and Reddit tasks and achieves the highest overall performance across categories, surpassing AgentOccam by $\sim$6% and outperforming BasicClaudeAction and WebArenaReplication by $\sim$20%. We further validate the generalizability of StructuredAgent using Kimi-K2 as the backbone model, with results presented in Figure˜2.
Method kimi Gemini GPT Avg
k2 ($\uparrow$) 2.5F ($\uparrow$) 4.1 ($\uparrow$)
StructuredAgent 56.7 36.7 46.7 46.7
StructuredAgentMem 55.0 36.7 45.0 45.6
AgentOccam 45.0 30.0 41.7 38.9
BasicClaudeAction 28.3 21.7 20.0 23.3
SteP 30.0 13.3 11.7 18.3
Table 3: Success rates (%) on Complex Shopping (Amazon Hard) tasks using Claude 3.7 Sonnet as the backbone, evaluated with three LLM judges: Kimi-k2, Gemini-2.5-Flash, and GPT-4.1. Bold indicates the best performance per column. StructuredAgent achieves the highest average at 46.7%, outperforming all baselines across every evaluator.
RQ2) Does StructuredAgent continue to benefit from hierarchical planning with a stronger model from the same family? Table˜3 compares agent performance on Amazon Hard tasks using Claude 3.7 as the backbone model. Despite Claude 3.7 being explicitly trained on agentic tasks, StructuredAgent still outperforms AgentOccam by 7.8% and BasicClaudeAction by 23.4%, demonstrating that the benefits of hierarchical planning persist even with stronger base models. We also observe that the Structured Memory module provides no additional gain when Claude 3.7 serves as the backbone. This suggests that the stronger model effectively leverages the And/Or tree to track constraints and maintain awareness of alternative solutions without requiring explicit structured memory support.
RQ3) Does StructuredAgent generalize to model families beyond Claude? Table˜4 evaluates StructuredAgent and AgentOccam using Kimi-k2-0905 as the backbone model on a subset of WebArena tasks that do not rely on string-based or URL-based evaluation metrics. StructuredAgent outperforms AgentOccam on 3 out of 4 task categories and achieves a higher overall success rate, demonstrating that the benefits of our framework generalize across model families. We attribute the relatively small performance margin compared to Claude-based results to the degraded performance of Kimi-k2 on long-context inputs, which is a known limitation of this model.
RQ3) How does StructuredAgent compare with other agents in terms of trajectory length and runtime efficiency? Fig. 6 presents violin plots of the number of browser actions that each agent takes per task on WebArena, and Fig. 9 shows the corresponding wall-clock time in minutes. The overlaid marker within each violin plot denotes the per-agent mean. StructuredAgent takes a moderately larger number of actions and more time on Map and GitLab tasks, which are categories where all agents achieve relatively lower success rates. This pattern suggests that StructuredAgent engages in increased exploration on harder tasks where straightforward strategies prove insufficient. On Reddit and Shopping tasks, where StructuredAgent achieves higher success rates, it takes a comparable number of actions to AgentOccam. BasicClaudeAction and SteP consistently generate the longest trajectories across most task categories, which aligns with their lower overall success rates.
Figure˜7 shows the analogous violin plot distributions for Amazon Easy and Amazon Hard tasks. StructuredAgent takes more actions than AgentOccam across both task types. The increase remains modest on easier tasks, staying less than twice the number of actions that AgentOccam takes, while StructuredAgent produces substantially longer trajectories on hard tasks. Qualitative inspection of these trajectories reveals that StructuredAgent systematically explores multiple solution strategies before abandoning a task, as the example in Table˜11 illustrates.
5 Conclusion
We combine dynamic And/Or tree-based planning with structured memory to tackle complex, long-horizon web tasks that challenge existing agentic LLMs. The And/Or tree decomposes tasks hierarchically, enabling principled error recovery and parallel exploration of alternative strategies, while the structured memory module tracks multi-constraint state across extended retrieval tasks. Evaluating on a complex shopping benchmark, WebArena, and WebVoyager, our framework consistently and significantly outperforms strong baselines across multiple model families, demonstrating robust and generalizable improvements in web-browsing agents.
Appendix A EXPERIMENT Results
Method Map Shop. Reddit GitLab Overall
(81) (126) (11) (63) (281)
StructuredAgent 0.348 0.432 0.545 0.383 0.401
AgentOccam 0.410 0.370 0.591 0.410 0.399
Table 4: Success rates on WebArena using Kimi-K2 as the backbone model, evaluated on the subset of WebArena tasks with URL and string evaluators.
(a) GITLAB
(b) MAP
(c) REDDIT
(d) SHOPPING
Figure 9: Completion time distributions (in minutes) for StructuredAgent and AgentOccam across four WebArena task categories, with the mean marked per violin. StructuredAgent exhibits longer completion times and heavier right tails, reflecting greater variance in task duration due to its deliberate planning and dynamic replanning upon encountering failures.
(a) Amazon Easy
(b) Amazon Hard
(c) WebVoyager Easy
Figure 10: Average action counts per task for each agent across Amazon Easy, Amazon Hard, and WebVoyager Easy, broken down by action type (click, type, note, go_back). StructuredAgent and StructuredAgentMem exhibit notably higher type and note counts on harder tasks, reflecting more thorough information gathering and constraint tracking.
(a) GITLAB
(b) MAP
(c) REDDIT
Figure 11: Average action counts per task for each agent across three WebArena task categories (GitLab, Map, Reddit), broken down by action type (click, type, note, go_back, go_home). StructuredAgent exhibits notably higher note counts, reflecting its structured information tracking, while maintaining competitive click and type counts relative to baselines.
Appendix B Tree and Agent Operations Details
Figure 12: Example of Structured Memory used to track candidate solutions during information-seeking tasks.
B.1 Planning and Execution Algorithm
Node Expansion: This operm ator identifies a node’s type and, when appropriate, generates its children. For nodes, it generates subgoal descriptions; for Action nodes, it produces a browser-level command; and for Or nodes, it outputs a set of candidate strategies along with their scores. The framework then updates the tree structure accordingly. If the expansion fails, the agent retries the operation up to a predefined limit. After repeated failures, the agent marks the node as Failed.
Node Repair: This operator revises or Or nodes that have entered the Failed state. Conditioned on the node’s metadata and contextual summaries, it proposes modifications to the node’s children. When the node remains repairable, the operator recommends adding new children; otherwise, it instructs the framework to prune the node. If several repair attempts fail, the agent prunes the node.
Global Tree Update: This operator updates the global tree structure based on newly observed information. Given the current tree and contextual metadata, it issues instructions for pruning irrelevant or infeasible subtrees and for updating node descriptions. This operator aggressively removes unpromising branches to maintain tractability. If repeated inference attempts fail, the operator recommends fallback pruning strategies.
Node Completion Check: This operator evaluates whether an node has satisfied its objective based on the statuses of its children. The framework invokes this operator once all children have completed execution and at least one has succeeded. The operator assesses the node’s description, the task progress summary, the notes summary, and the task constraints, and returns a binary decision indicating whether the node’s objective has been achieved.
Observation Summarizer: This module that maintains the agent’s internal representations and tracks its progress. This operator processes new observations to update the task progress summary, generate an observation summary, and provide guidance for subsequent actions. When the agent encounters a new webpage, this module extracts task-relevant information, updates the agent’s notes, and produces a revised notes summary. These summaries enable the agent to maintain an accurate understanding of the environment and support effective planning.
Table 5: Internal Representations Maintained by StructuredAgent
Representation Description
Interaction History Logs all observation summaries and actions up to the current time step
Task Progress Summary Tracks progress toward the main objective
Notes Summary Aggregates information from all past observations that are relevant to the task
Task Constraints Encodes global and item-level requirements from the task description
Observation Summary Distills salient information from the current web page at each step
Action History Lists all executed browser actions
Appendix C Related Works
Prior work on web agents can be broadly be classified into two main categories: (1) LLM inference-based web agents and (2) fine-tuned LLMs as web agents.
LLM Inference-Based Web Agents:
Several prior works [Putta et al., 2025, Zhang et al., 2024, He et al., 2024, Zhou et al., 2024] leverage closed-source large language or multimodal models for complex web-based tasks, often by employing inference-time strategies. For example, Wang et al. [2025] use LLMs to extract patterns from examples and prior trajectories to guide navigation and decision-making. Other approaches incorporate Monte Carlo Tree Search (MCTS) [Zhang et al., 2024] or value function networks [Koh et al., 2024] to facilitate exploration and backtracking during planning. Methods such as Reflexion [Shinn et al., 2023] add reflection mechanisms and evaluators, though often at the cost of frequent resets.
Hierarchical planning has also been investigated as a means of addressing compositional web tasks. Erdogan et al. [2025] introduce an optimization layer for higher-level planning, while Sodhi et al. [2024] propose dynamic multi-level control, though their approach depends on domain-specific prompt engineering and task-specific policies. Yang et al. [2025] present a lightweight tree structure with effective text filtering and incremental plan revision, which outperforms many baselines but struggles on more complex tasks due to limited error recovery and weak information retention. In contrast, our method develops a more general And/Or planning tree that supports robust error handling and strategic planning with explicit dependency tracking, yielding more interpretable trajectories (see Table 10).
Fine-tuned LLMs as Web Agents:
Recent work has investigated the use of open-source language models for web-based tasks, though these models often struggle with high failure rates attributable to their relatively modest scale and limited domain-specific knowledge. To bridge this gap, subsequent research proposed fine-tuning open-source models on expert trajectories sourced from either human demonstrators or proprietary LLMs, demonstrating that such fine-tuning yields meaningful performance gains, particularly when evaluation tasks are drawn from the same distribution as the training data.
The training paradigms employed in this line of work generally fall into two categories: imitation learning, in which agents are trained offline on pre-collected expert demonstrations, and reinforcement learning (RL), in which agents improve through direct interaction with the environment. Imitation learning approaches train agents to maximize the likelihood of expert actions, but this objective makes them brittle and poorly suited to complex or out-of-distribution scenarios. Their success is heavily contingent on the quality and breadth of the demonstration data, which is costly to obtain, especially when relying on proprietary models for trajectory generation.
RL-based training, while offering a path toward more adaptive agents, introduces its own set of challenges. Online RL for web agents is computationally demanding, and agents typically receive only sparse binary reward signals tied to overall task completion. This sparsity gives rise to a credit assignment problem: when a task fails, the learning algorithm may inadvertently penalize intermediate actions that were, in fact, correct. Consequently, fine-tuning capable and generalizable web agents using high-quality trajectories paired with dense reward signals remains a largely underexplored direction in the field.
Figure 13: Human Decomposition Intervention in AND/OR Tree Planning to Correct Bad Subtask Decomposition. The diagram illustrates how a human decomposition intervention can redirect an AI agent that has produced an incorrect subgoal breakdown for a multi-step task. The task requires finding the driving route between the home cities of the 1980 and 1991 Super Bowl champions. Without intervention (bottom), the agent incorrectly assumes destinations a priori, defaulting to Pittsburgh, PA and Washington, DC, without first verifying which teams won in those respective years, resulting in a hallucinated but confidently stated incorrect answer. With intervention (top), a human intercepts the agent’s flawed decomposition and injects three corrective subgoals into the AND/OR tree: (1) identify the 1980 Super Bowl champions and their home city, (2) identify the 1991 Super Bowl champions and their home city, and (3) return to the mapping tool with the verified locations. This human-guided correction grounds the plan in factual information before execution continues, yielding the correct route from Pittsburgh, PA (Pittsburgh Steelers) to East Rutherford, NJ (New York Giants) at 599 km and approximately 7 hours 23 minutes. For simplicity, we illustrate the human intervention at the first level of the AND/OR planning tree; however, such human decomposition interventions can be applied at any level of the tree, allowing a human operator to catch and correct errors in subgoal formulation at whatever level of abstraction they arise.
Appendix D Algorithm Details
D.1 Planning and Execution Algorithm
The agent begins by pushing the root node which represents the given task, onto the DFS stack in the ENTERING state. As in classical iterative DFS, the algorithm then iteratively pops the top node from the stack and processes it according to its current state.
When processing a node in the Entering state, the algorithm first checks the node’s status. By default, new nodes are marked as unvisited. If the node is encountered for the first time, the algorithm invokes the Node Expansion module to determine its type and expand it. For nodes, the module generates an ordered list of children according to task constraints. For Or nodes, it ranks alternatives by their estimated likelihood of success, as determined by LLM scoring or a value function. For Action nodes, it outputs the precise browser command to execute. After expansion, or if the node was previously expanded, the agent marks the node as visited and re-adds it to the stack in the EXITING state.
Next, the agent proceeds according to node type. For nodes, it pushes each child onto the stack in the Entering state for further processing. For Or nodes, it pushes only the most promising unprocessed child in the Entering state. If the node has no children, no further action is taken until its status is evaluated in the Exiting state. For Action nodes, the agent immediately executes the specified browser command. If the command is successful, the agent marks the node as successful, updates its internal context via the Observation and Notes Summarization module. The Global Tree Update module prunes irrelevant branches and refines node descriptions as needed, propagating these changes upward to keep the stack and tree consistent. If the Action node fails, the agent marks it as FAILED.
When processing a node in the Exiting state, the agent evaluates completion criteria based on the node type. For nodes, the agent verifies that all valid, that is, unpruned or undeleted, child nodes have succeeded and invokes the Node Completion module to ensure that the overall objective of the node is met. Although this check is theoretically applicable to all nodes, we found it to be overly strict in practice. Therefore, to avoid unnecessary retries and failures, we apply this strict completion check only at the root node in our implementation. If completion is verified, the agent marks the node as successful, otherwise, as FAILED. For Or nodes, the agent checks if at least one child has succeeded, marking the node as SUCCESS if so, and as FAILED otherwise. For Action nodes, the agent sets the status to success or FAILED depending on the outcome of the browser operation. Once a node is marked as success, it is not processed again.
If the agent processes a node in the Failed state, it attempts to repair the node or marks it as pruned. For Action nodes, it prunes the node and updates the stack. If a failed Action node belongs to an node, the agent deletes all remaining unexecuted siblings, since the conjunctive requirement cannot be satisfied. For nodes, if none of the children have achieved the node’s objective, the agent checks the node’s revision count. If the revision limit is not reached, the Node Repair module attempts to generate new subgoals or modify the set of children. If repair fails or the revision limit is reached, the agent prunes the node and all its descendants, propagating failure upward and updating the stack. For Or nodes in the Failed state, the agent first checks for any remaining unprocessed children. If such children exist, it selects and pushes the most promising one onto the stack in the Entering state. If all alternatives are exhausted, it attempts node repair if within the revision limit, otherwise, it prunes the node. As with nodes, any pruning or failure of an Or node propagates upward, ensuring the global plan remains consistent and that failures at lower levels trigger repair or pruning at higher levels.
This DFS-based algorithm allows the agent to incrementally build and adapt its hierarchical plan, efficiently responding to new observations and failures, while executing the task.
// Perform rollback when parent is an OR node
if node.parent and node.parent.type == NodeType.OR then
  success $\leftarrow$ PERFORM_ROLLBACK(node.parent)
 if success == true then
     currentUrl $\leftarrow$ node.parent.url
  end if
end if
// Increase node execution count
node.execution_count $\leftarrow$ node.execution_count + 1
// If node hasn’t yet been visited, expand it
if node.type == NodeType.UNKNOWN and node.status $\notin$ CLOSED_STATUSES then
  // Populate the node type
  node, _, ACTION $\leftarrow$ POPULATE_NODE_TYPE(node)
end if
node.status $\leftarrow$ NodeStatus.VISITED
// Process node based on its type
if node.type == NodeType.ACTION then
  stack, node $\leftarrow$ PROCESS_ACTION_NODE(node, stack)
  // Algorithm 6
end if
else if node.type == NodeType.AND then
  stack, node $\leftarrow$ PROCESS_AND_NODE(node, stack)
  // Algorithm 7
end if
else if node.type == NodeType.OR then
  stack, node $\leftarrow$ PROCESS_OR_NODE(node, stack)
  // Algorithm 8
end if
return stack, node
Algorithm 2 PROCESS_NODE_ENTERING(node, stack)
// ACTION NODE IN EXITING STATE
if node.type == NodeType.ACTION then
  // Put them back into the stack in failure mode
  if node.status $\in$ FAILED_OR_PRUNED then
     stack.append((node, NodeState.FAILED))
  end if
 else
     node.status $\leftarrow$ NodeStatus.SUCCESS
  end if
end if
// ACTION NODE IN FAILED STATE
if node.type == NodeType.ACTION then
  node.status $\leftarrow$ NodeStatus.PRUNED
  BACKTRACK_FAILURE(node)
end if
// Process node based on its type
else if node.type == NodeType.AND then
  stack, node $\leftarrow$ PROCESS_FAILED_AND_NODE(node, stack)
  // Algorithm 9
end if
else if node.type == NodeType.OR then
  stack, node $\leftarrow$ PROCESS_FAILED_OR_NODE(node, stack)
  // Algorithm 10
end if
return stack, node
Algorithm 4 PROCESS_NODE_FAILED(node, stack)
// AND NODE IN EXITING STATE
else if node.type == NodeType.AND then
  success $\leftarrow$ IS_SUCCESSFUL_AND(node)
 if success == true then
     // If node revision has not reached max limit, check if node is complete
     // If not complete, repair the node
     if (node.revision_count << MAX_REVISION_COUNT) then
        isComplete, reasoning $\leftarrow$ CHECK_FOR_COMPLETION_AND(node)
        node.reasoning $\leftarrow$ reasoning
       if isComplete == true then
           node.status $\leftarrow$ NodeStatus.SUCCESS
           continue
        end if
       else
           // Is complete check failed, mark node as failed and push it on the stack in FAIL state
           node.status $\leftarrow$ NodeStatus.FAIL
           node.reasoning $\leftarrow$ reasoning
           stack.append((node, NodeState.FAILED))
           continue
        end if
     end if
    else
        // If node has already been revised sufficient number of times and the children are successful, mark the node as successful
        node.status $\leftarrow$ NodeStatus.SUCCESS
        continue
     end if
  end if
 else
     // Mark the node as failed and push it onto the stack in FAIL state
     node.status $\leftarrow$ NodeStatus.FAIL
     stack.append((node, NodeState.FAILED))
  end if
end if
// OR NODE IN EXITING STATE
else if node.type == NodeType.OR then
  success $\leftarrow$ IS_SUCCESSFUL_OR(node)
 if success == true then
     node.status $\leftarrow$ NodeStatus.SUCCESS
  end if
 else
     node.status $\leftarrow$ NodeStatus.FAIL
     stack.append((node, NodeState.FAILED))
  end if
end if
return stack, node
Algorithm 3 PROCESS_NODE_EXITING(node, stack)
if node.parent == null then
  return
  // No parent to backtrack to
end if
if node.parent.type == NodeType.AND and node.parent.ordered == true then
  // Handle ordered AND node: delete remaining siblings
  remaining $\leftarrow$ GET_REMAINING_EXCLUDING_NODE(node, node.parent)
  undeletedChildren $\leftarrow$ []
  foreach rem in remaining do
     if rem.status != NodeStatus.DELETED then
        undeletedChildren.append(rem)
        // Filter undeleted children
     end if
  end foreach
 if undeletedChildren not empty then
     foreach rem in undeletedChildren do
        rem.status $\leftarrow$ NodeStatus.DELETED
        // Mark as deleted
     end foreach
    deletedChildren $\leftarrow$ RECURSIVELY_DELETE_CHILDREN(undeletedChildren)
     if stack not empty then
        stack $\leftarrow$ PURGE_STACK(deletedChildren, stack)
        // Remove from stack
     end if
  end if
end if
Algorithm 5 BACKTRACK_FAILURE(node)
// If node represents an ACTION, execute it. If successful, globally update tree
status $\leftarrow$ null
// Before processing any node, push it back onto stack in EXITING state
stack.append((node, NodeState.EXITING))
success $\leftarrow$ false
// Execute action with retry logic
while node.retry_count << MAX_RETRY_COUNT do
  // Try block for attempting to perform action
  if NO_EXCEPTION then
     action, status $\leftarrow$ PERFORM_ACTION(node)
    if env.steps >> MAX_STEPS then
        return stack, node
        // Reached maximum steps
     end if
    if env.done() then
        return stack, node
        // Environment done after performing action
     end if
    // Check if action was successfully executed
     assert(status is True)
    // Extract notes from the action if there are any
     UPDATE_NOTES(node.action)
    // Mark ACTION node as successful
     node.status $\leftarrow$ NodeStatus.SUCCESS
     success $\leftarrow$ true
    // If ACTION has resulted in change inn current observation
     isNoteAction $\leftarrow$ IS_NOTE_ACTION(node.action)
    if not isNoteAction then
        // Update notes and observation summaries
        ASYNC_FULL_UPDATE()
        STEP $\leftarrow$ STEP + 1
       // Update URL of current node and current URL
        currentUrl $\leftarrow$ env.get_url()
        UPDATE_URL(node)
     end if
    node.retry_count $\leftarrow$ node.retry_count + 1
     break
  end if
 // If exception occurred, update retry count and continue
  node.retry_count $\leftarrow$ node.retry_count + 1
end while
// If success is false, update status to fail
if success == false then
  node.status $\leftarrow$ NodeStatus.FAIL
end if
return stack, node
Algorithm 6 PROCESS_ACTION_NODE(node, stack)
// For an AND node, process all unprocessed children and put them on stack
// First push the node back onto stack in EXITING state
// so that control returns to it once children have been executed
stack.append((node, NodeState.EXITING))
children $\leftarrow$ node.children
// Check if the AND node is valid
isValid $\leftarrow$ IS_VALID_AND(children)
// Check if the AND node is successful
success $\leftarrow$ IS_SUCCESSFUL_AND(node)
// If successful, mark node as successful
if success then
  node.status $\leftarrow$ NodeStatus.SUCCESS
end if
else if not isValid then
  // I there are no valid children, mark it as fail
  node.status $\leftarrow$ NodeStatus.FAIL
end if
else
  // Put unprocessed children in stack in reversed order
  foreach child in reverse(node.children) do
     if child.status $\notin$ CLOSED_STATUSES then
        stack.append((child, NodeState.ENTERING))
     end if
  end foreach
end if
return stack, node
Algorithm 7 PROCESS_AND_NODE(node, stack)
// For an OR node, process next promising child
// First push node back onto stack in EXITING state
// to return control to it after child has been processed
stack.append((node, NodeState.EXITING))
children $\leftarrow$ node.children
// Check if node is successful
success $\leftarrow$ IS_SUCCESSFUL_OR(node)
// There should be one valid child
isValid $\leftarrow$ IS_VALID_OR(children)
// Mark the node as successful
if success then
  node.status $\leftarrow$ NodeStatus.SUCCESS
end if
// If no children are available to process, mark node as failed
else if not isValid then
  node.status $\leftarrow$ NodeStatus.FAIL
end if
else
  // Find next promising child (children ordered by estimated likelihood of success)
  child $\leftarrow$ FIND_NEXT_PROMISING(children)
 // Push promising child onto the stack
  stack.append((child, NodeState.ENTERING))
end if
return stack, node
Algorithm 8 PROCESS_OR_NODE(node, stack)
// For an AND node in FAILED state, revise, prune, or backtrack
// Initialize tracking variables
delChildren $\leftarrow$ []
addedChildren $\leftarrow$ []
prunedChildren $\leftarrow$ []
revised $\leftarrow$ false
isValid $\leftarrow$ IS_VALID_AND(node.children)
// If node is invalid but has at least one successful child, check if the node’s objective has been completed
if not isValid then
  reasoning $\leftarrow$ ""
  hasOneSuccess $\leftarrow$ HAS_ATLEAST_ONE_SUCCESS(node)
 if hasOneSuccess then
     isComplete, reasoning $\leftarrow$ CHECK_FOR_COMPLETION_AND(node)
     node.reasoning $\leftarrow$ reasoning
    if isComplete and isComplete == true then
        node.status $\leftarrow$ NodeStatus.SUCCESS
        return stack, node
     end if
  end if
 // If node revision has not reached the maximum limit, repair it
  node.status $\leftarrow$ NodeStatus.FAIL
 if (len(node.children) << MAX_CHILDREN and node.revision_count << MAX_REVISION_COUNT) then
     // Revise the node
     node, delChildren, prunedChildren, addedChildren $\leftarrow$ REVISE_AND_NODE(node)
    removedChildren $\leftarrow$ delChildren + prunedChildren
    // Update stack and purge all the removed children
     stack $\leftarrow$ PURGE_STACK(removedChildren, stack)
    // Check if node has been revised. If revised, update revision count
     if node.status != NodeStatus.PRUNED and (len(addedChildren) != 0 or len(prunedChildren) != 0) then
        revised $\leftarrow$ true
        node.revision_count $\leftarrow$ node.revision_count + 1
     end if
  end if
end if
// If node is successful, mark the node as successful
if node.status == NodeStatus.SUCCESS then
  RECURSIVELY_MARK_SUCCESS(node.id)
end if
// If the node has been revised or there are children to be processed, then push the node back onto the stack in ENTERING state
else if revised or isValid then
  node.status $\leftarrow$ NodeStatus.VISITED
  stack.append((node, NodeState.ENTERING))
end if
// If node is invalid and not revised, prune the node
else if not revised and not isValid then
  prunedNodes $\leftarrow$ RECURSIVELY_PRUNE(node.id)
 // Update stack and purge all the pruned nodes
  stack $\leftarrow$ PURGE_STACK(prunedNodes, stack)
 // Propagate failure
  BACKTRACK_FAILURE(node)
end if
return stack, node
Algorithm 9 PROCESS_FAILED_AND_NODE(node, stack)
// If an OR node is in failed state, either execute the next promising child
// (if any) or repair or prune the node. If node is pruned and the node’s parent is AND,
// then purge the rest of the children, or else if the parent is OR, then process the next promising child
revised $\leftarrow$ false
addedChildren $\leftarrow$ []
isValid $\leftarrow$ IS_VALID_OR(node.children)
if isValid then
  node.status $\leftarrow$ NodeStatus.VISITED
  stack.append((node, NodeState.ENTERING))
  return stack, node
end if
else
  // If the node revision count has not reached the max limit, revise the node
  if node.revision_count << MAX_REVISION_COUNT then
     // Revise OR node
     node, addedChildren, prunedChildren $\leftarrow$ REVISE_OR_NODE(node)
    // Update stack
     stack $\leftarrow$ PURGE_STACK(prunedChildren, stack)
    // Check if node was revised, update revision count
     if len(addedChildren) >> 0 then
        revised $\leftarrow$ true
        node.revision_count $\leftarrow$ node.revision_count + 1
     end if
  end if
 // If node was revised, push it back onto the stack in ENTERING STATE
  if revised then
     node.status $\leftarrow$ NodeStatus.VISITED
     // If new children were added, put this node back onto the stack
     stack.append((node, NodeState.ENTERING))
  end if
 else
     // If not revised then prune the node
     prunedNodes $\leftarrow$ RECURSIVELY_PRUNE(node.id)
    // Update stack
     stack $\leftarrow$ PURGE_STACK(prunedNodes, stack)
    // Propagate failure and process next promising child of OR node
     BACKTRACK_FAILURE(node)
  end if
end if
return stack, node
Algorithm 10 PROCESS_FAILED_OR_NODE(node, stack)
Internal Functions Reference
Table 6: Node Status Definitions in StructuredAgent Framework
Status Description
UNVISITED Initial state of all nodes before processing begins.
VISITED Node has been processed but final outcome is not yet determined.
SUCCESS Node’s objective is achieved. For nodes, all children succeed and collectively satisfy the objective; for Or nodes, any single child’s success suffices; for Action nodes, the corresponding browser operation succeeds.
FAIL Node’s objective cannot be achieved. For nodes, occurs if any child fails or is pruned; for Or nodes, if the selected child fails; for Action nodes, if the operation is invalid or unsuccessful.
PRUNED Node becomes irrelevant or irresolvable due to new information discovered during execution.
DELETED Node is marked as inapplicable due to failure or pruning of a predecessor sibling in an node with ordering constraints.
Function Name Description
PERFORM_ROLLBACK(node) Performs context rollback to restore the browser state to the URL of the given node. Returns success status. Used when processing OR node children to ensure clean state.
POPULATE_NODE_TYPE(node) Expands an UNKNOWN node by determining its type (ACTION, AND, OR) and populating relevant fields like children or action details. Returns updated node, status, and action.
BACKTRACK_FAILURE(node) Handles backtracking when a node fails. For ordered AND nodes, deletes remaining siblings and their descendants. Propagates failure up the tree.
PERFORM_ACTION(node) Executes the action specified in the node (e.g., click, type, navigate). Interacts with the environment to perform the web automation task. Returns action and execution status.
UPDATE_NOTES(action) Extracts and updates notes from the action if any are present. Notes are used to record information during execution.
IS_NOTE_ACTION(action) Checks if an action is a "note" action. Note actions don’t change observations, so they don’t require a full tree update.
ASYNC_FULL_UPDATE() Performs a global asynchronous update of the agent’s state, observation summaries, and tree structure after an observation-changing action.
UPDATE_URL(node) Updates the URL field of the given node to match the current environment URL.
IS_VALID_AND(children) Checks if an AND node is valid by verifying it has at least one non-closed child that can be processed.
IS_SUCCESSFUL_AND(node) Checks if an AND node is successful by verifying all its children have succeeded.
CHECK_FOR_COMPLETION_AND(node) Evaluates whether an AND node’s objective has been completed even if not all children succeeded. Returns completion status and reasoning. Used to determine if partial completion is acceptable.
HAS_ATLEAST_ONE_SUCCESS(node) Checks if a node has at least one successful child. Used to determine if completion checking should be performed for AND nodes.
IS_SUCCESSFUL_OR(node) Checks if an OR node is successful by verifying at least one of its children has succeeded.
IS_VALID_OR(children) Checks if an OR node is valid by verifying it has at least one child available to process.
FIND_NEXT_PROMISING(children) Finds the next most promising child to explore from a list of OR node children. Children are ordered by estimated likelihood of success.
REVISE_AND_NODE(node) Revises a failed AND node by analyzing failures, potentially pruning unsuccessful children, and adding new children. Returns updated node, deleted children, pruned children, and added children.
REVISE_OR_NODE(node) Revises a failed OR node by generating new alternative approaches. Returns updated node, added children, and pruned children.
GET_REMAINING_EXCLUDING_NODE(node, parent) Returns all sibling nodes that come after the given node in the parent’s children list. Used for deleting remaining children in ordered AND nodes.
RECURSIVELY_DELETE_CHILDREN(children) Marks the given children and all their descendants as deleted in the tree structure. Returns list of deleted nodes.
RECURSIVELY_PRUNE(node_id) Recursively prunes a node and all its descendants, marking them as pruned in the tree structure. Returns list of pruned nodes.
RECURSIVELY_MARK_SUCCESS(node_id) Recursively marks a node and its relevant descendants as successful. Used to propagate success status through the tree.
PURGE_STACK(deletedChildren, stack) Removes all deleted or pruned nodes from the execution stack to prevent processing invalid branches. Returns updated stack.
Node Status and Type Collections
Collection Description
CLOSED_STATUSES Node statuses that indicate the node is no longer active or executable: [SUCCESS, PRUNED, DELETED]
FAILED_STATUSES Node statuses that imply failure or deleted: [PRUNED, DELETED]
FAILED_OR_PRUNED Node statuses that indicate failure or pruned: [FAIL, PRUNED]
VISITED_TYPES Node statuses that indicate a node has been visited or processed in some way: [FAIL, VISITED, PRUNED, SUCCESS]
VALID_NODE_TYPES All valid/expected node types (excluding UNKNOWN): [AND, OR, ACTION]
NON_LEAF_NODES Node types that represent composite (non-leaf) logic structures: [AND, OR]
Table 9: And/OrTree Node Attributes
Attribute Type Description
type NodeType The type of the node (e.g., AND, OR, ACTION)
description str A human-readable description of the node
id str A unique identifier for the node
steps int Number of steps or operations this node represents (default: 0)
children List[Node] A list of child nodes (default: empty list)
parent Optional[Node] A reference to the parent node, if any
status NodeStatus The current execution status of the node (default: UNVISITED)
execution_count int Number of times this node has been executed (default: 0)
retry_count int Number of retry attempts for this node (default: 0)
revision_count int Number of revisions or changes to this node (default: 0)
depth int Depth of the node within the tree (root is 0, default: 0)
ordered bool Whether the execution of child nodes is ordered (default: True)
url str Optional URL associated with the node (default: empty string)
action str Action or command associated with the node (default: empty string)
reasoning str Explanation or rationale behind this node’s inclusion or action (default: empty string)
others dict Arbitrary additional metadata or context (default: empty dict)
notes List[str] Freeform notes or comments related to the node (default: empty list)
Appendix E Planning Tree
Figure 14: AgentOccam vs StructuredAgent on a Pinterest task: AgentOccam terminates its search prematurely when it fails to explore the third trend. On the other hand, StructuredAgent is able to successfully retrieve items for all three trending women’s ethnic wear styles.
- [0] (AND) Research Pinterest to identify the top 3 women’s ethnic wear trends for Diwali. For each trend: 1) Specify the signature color currently trending. 2) Find one matching footwear style and one purse, both in the same color (same color as the trend) from the Pinterest Shop. 3) Provide the name of a potential brand or seller for each footwear and purse, and include a direct source or pin link if possible. Focus on vibrant, festive, and popular looks. Do not end your search early. Limit your review to only the first 2 scrolls of results. Do not click on external links. [success] - [0.1] (Atomic) Search for Diwali ethnic wear trends on Pinterest (type [401] [Diwali ethnic wear trends 2024] [1]) [success] - [0.2] (Atomic) Document 3 trending ethnic wear styles with their signature elements (note [Trending Diwali Ethnic Wear 2024: 1. Pink - Anupa Sari with embellished blouse from INDIASPOPUP.COM 2. Pista Green - Crush fabric sari with sequins embroidery border and yellow jacquard blouse from Etsy 3. Powder Peach - Stripe sari set with thread, pearl and sequin embroidery, featuring plunging neckline blouse from KYNAH]) [success] - [0.3] (AND) Navigate to Pinterest Shop and search for footwear in Pink, Pista Green, and Powder Peach shades from verified sellers specifically for Diwali looks [success] - [0.3.1] (Atomic) Navigate to Pinterest Shop section (click [208]) [success] - [0.3.2] (Atomic) type [401] [Pink ethnic juttis Diwali footwear] [1] (type [401] [Pink ethnic juttis Diwali footwear] [1]) [success] - [0.3.3] (Atomic) Return to previous page after completing pink footwear search (go_back) [success] - [0.3.4] (Atomic) type [401] [Pista Green ethnic juttis Diwali] [1] (type [401] [Pista Green ethnic juttis Diwali] [1]) [success] - [0.3.5] (Atomic) Return to previous page to continue accessory search (go_back) [success] - [0.3.6] (Atomic) type [401] [Powder Peach ethnic juttis Diwali] [1] (type [401] [Powder Peach ethnic juttis Diwali] [1]) [success] - [0.4] (OR) Search Pinterest Shop for matching ethnic clutches/purses in three colors by trying alternative search terms: 1) Search "Pink potli Diwali bag" or "Pink ethnic clutch" from verified sellers, 2) Search "Pista Green potli Diwali bag" or "Mint ethnic clutch" from verified sellers, 3) Search "Peach potli Diwali bag" or "Blush ethnic clutch" from verified sellers. Limit to first 2 scrolls of results. [success] - [0.4.1] (AND) Search for powder peach ethnic clutches in Shop section [success] - [0.4.1.1] (Atomic) Search for powder peach ethnic clutches in Shop section (type [401] [powder peach ethnic clutch Diwali] [1]) [success] - [0.4.1.2] (Atomic) Execute search query for powder peach ethnic clutches (type [401] [powder peach ethnic potli clutch bag Diwali] [1]) [success] - [0.4.1.4] (Atomic) note [Found matching powder peach potli bags from verified sellers: 1. 5 Elements Peach Brocade Pattern Potli Bag at Pernia’s Pop Up Shop 2. AMYRA Peach Poly Silk Pearl Embroidered Potli Bag 3. SWABHIMANN ACCESSORIES Peach Pure Silk Potli Bag] (note [Found matching powder peach potli bags from verified sellers: 1. 5 Elements Peach Brocade Pattern Potli Bag at Pernia’s Pop Up Shop 2. AMYRA Peach Poly Silk Pearl Embroidered Potli Bag 3. SWABHIMANN ACCESSORIES Peach Pure Silk Potli Bag]) [success] - [0.4.1.5] (Atomic) Modify search query to find matching powder peach potli bags (type [401] [powder peach potli bag Diwali] [1]) [success] - [0.4.1.6] (Atomic) click [208] to ensure Shop filter is active before searching for powder peach accessories (click [208] to ensure Shop filter is active before searching for powder peach accessories) [success] - [0.4.1.8] (Atomic) Modify search to find powder peach potli bags for Diwali (type [401] [powder peach potli bag Diwali] [1]) [success] - [0.5] (Atomic) Document powder peach accessory pairs with seller details (note [Powder Peach - Footwear: Coral Haze Juttis, Jutti Express formal wear, 5 Elements ethnic shoes; Purse: 5 Elements Brocade Potli with tassels, AMYRA Poly silk Pearl Embroidered Potli, Ri Ritu Kumar Net and Silk Zari Potli]) [success] - [0.6] (AND) Document pink and pista green accessory pairs with seller details and pin links, ensuring color coordination and authenticity of sellers [success] - [0.6.1] (AND) Search and document pink ethnic footwear with seller details [success] - [0.6.1.1] (Atomic) Search for pink ethnic footwear for Diwali (type [401] [pink ethnic juttis Diwali] [1]) [success] - [0.6.1.2] (Atomic) note [Found matching pink footwear options: 1. B Anu Designs - Baby Pink Punjabi Jutti with Mirror and Bead handwork 2. Miar Designs - Pink Leather Embroidered Juttis at Pernia’s Pop Up Shop 3. YASSIO - Pink Silk Floral Embroidered Juttis] (note [Found matching pink footwear options: 1. B Anu Designs - Baby Pink Punjabi Jutti with Mirror and Bead handwork 2. Miar Designs - Pink Leather Embroidered Juttis at Pernia’s Pop Up Shop 3. YASSIO - Pink Silk Floral Embroidered Juttis]) [success] - [0.6.1.3] (Atomic) Return to previous page to continue accessory search (go_back) [success] - [0.6.2] (Atomic) Document matching pink potli/clutch from search results (note [Found matching pink brocade potli bag from 5 Elements with tassels and beads embellishments, suitable for festive wear]) [success] - [0.6.3] (Atomic) Return to previous page after powder peach potli search (go_back) [success] - [0.6.4] (Atomic) Search for pista green ethnic footwear using specific search terms (type [401] [Pista green embroidered juttis Diwali] [1]) [success] - [0.6.5] (Atomic) Search for matching pista green clutch/potli (type [401] [Pista green brocade potli Diwali] [1]) [success] - [0.6.6] (Atomic) Document complete accessory details with seller information for pista green trend accessories: Footwear - House of Vian’s Crepe Embroidered Juttis and 5 Elements’ Leatherette Embellished Juttis; Purse - 5 Elements Green Brocade Embroidered Potli with tassels and dupion lining, and Tarini Nirula Green Brocade Stone Embellished Potli with antique gold frame (Document complete accessory details with seller information for pista green trend accessories: Footwear - House of Vian’s Crepe Embroidered Juttis and 5 Elements’ Leatherette Embellished Juttis; Purse - 5 Elements Green Brocade Embroidered Potli with tassels and dupion lining, and Tarini Nirula Green Brocade Stone Embellished Potli with antique gold frame) [success] - [0.6.7] (Atomic) Search for powder peach ethnic footwear (type [401] [powder peach ethnic juttis Diwali] [1]) [success] - [0.6.10] (AND) Search for pink ethnic footwear pins from verified sellers and save direct links [success] - [0.6.10.1] (Atomic) View details of pink leather embroidered juttis for Pink category (click [17999]) [success] - [0.6.10.2] (Atomic) note [Pink leather juttis by Miar Designs - Features: Leather base with coin, thread, pearl, sequins and beads embroidery, Available at Pernia’s Pop Up Shop, Price: 40,00, Formal occasion wear] (note [Pink leather juttis by Miar Designs - Features: Leather base with coin, thread, pearl, sequins and beads embroidery, Available at Pernia’s Pop Up Shop, Price: 40,00, Formal occasion wear]) [success] - [0.6.10.3] (Atomic) Return to previous page to continue purse search (go_back) [success] - [0.6.11] (Atomic) Search for matching pink potli/clutch pins (type [401] [pink ethnic potli clutch Diwali] [1]) [success] - [0.6.14] (Atomic) Verify and document color coordination between footwear-purse pairs (note [Color coordination verified: Pink category - Pink Beaded Mirror Work Potli and Pink Brocade Potli perfectly match the pink leather juttis from Miar Designs in shade and festive style]) [success] - [0.7] (Atomic) Document direct pin links for pink accessories from verified sellers (note [Found verified seller pin links: 1. Pink Brocade Tassels Embellished Potli by 5 Elements - Direct Pin ID 20730 2. Pink Brocade Ghungroo Embroidered Potli by 5 Elements - Direct Pin ID 20752 3. Pink Beaded Mirror Work Potli by Etsy verified seller - Direct Pin ID 21242]) [success] - [0.8] (AND) Review and document direct pin links for all three color categories ensuring each has verified footwear and purse pairs with seller details [success] - [0.8.1] (Atomic) Document completed findings for Pink category (note [Pink Category Complete - Footwear: Miar Designs, Purses: Pink Beaded Mirror Work Potli (Etsy), Pink Brocade Potli (5 Elements)]) [success] - [0.8.2] (Atomic) Navigate to Pinterest Shop section for ethnic potli clutches (click [20411]) [success] - [0.8.7] (Atomic) Search for powder peach potli options (type [401] [powder peach ethnic potli clutch Diwali] [1]) [success] - [0.8.8] (Atomic) Search for matching Pista Green potli options with verified sellers and clear product details (type [401] [pista green ethnic potli clutch Diwali] [1]) (Search for matching Pista Green potli options with verified sellers and clear product details (type [401] [pista green ethnic potli clutch Diwali] [1])) [success] - [0.8.9] (Atomic) Document verified seller information for Pista Green potlis (note [Verified Pista Green Potlis: 1. PLODE - Green Velvet Swarovski Triangle Potli with crystal embellishments and rhinestone strap 2. Velleso - Green Mirror Work Potli with crystal drops and silver threadwork (24x23cm) 3. The Pink Potli - Green Raw Silk Potli with bead work and tassel embellishments]) [success] - [0.8.10] (Atomic) Search for matching Powder Peach ethnic footwear options (type [401] [powder peach ethnic footwear Diwali] [1]) [success] - [0.8.11] (Atomic) Search for matching Pista Green ethnic footwear options with verified sellers and clear product details (type [401] [pista green ethnic footwear Diwali] [1]) [success] - [0.8.12] (Atomic) Document pin links and details for Pista Green category footwear and purse pairs (note [Found direct pin links for Pista Green category: 1) House of Vian Green Crepe Embroidered Juttis with crystals and gota patti from Pernia’s Pop Up Shop 2025, 2) 5 Elements Green Leatherette Embellished Juttis with beadwork from Pernia’s Pop Up Shop 2025]) [success]
Table 10: Planning tree for the task “Research Pinterest to identify the top 3 women’s ethnic wear trends for Diwali. For each trend: 1) Specify the signature color currently trending. 2) Find one matching footwear style and one purse, both in the same color (same color as the trend) from the Pinterest Shop. 3) Provide the name of a potential brand or seller for each footwear and purse, and include a direct source or pin link if possible. Focus on vibrant, festive, and popular looks. Do not end your search early. Limit your review to only the first 2 scrolls of results. Do not click on external links.”
Table 11: Trajectory and final response comparison on task Diwali-05: “Research Pinterest to identify the top 3 women’s ethnic wear trends for Diwali. For each trend: 1) Specify the signature color currently trending. 2) Find one matching footwear style and one purse, both in the same color (same color as the trend) from the Pinterest Shop. 3) Provide the name of a potential brand or seller for each footwear and purse, and include a direct source or pin link if possible. Focus on vibrant, festive, and popular looks. Do not end your search early. Limit your review to only the first 2 scrolls of results. Do not click on external links.” Both agents use claude-3-5-sonnet-20241022.
Step StructuredAgent AgentOccam
1 search “Diwali ethnic wear trends 2024” search “Diwali ethnic wear trends 2024”
2 click search bar click first result
3 search “light pink ethnic purse palazzo style” click search bar
4 click product result [note] Top 3 trends: (1) Palazzo suits – Light Pink, (2) Stripe sari sets – Powder Peach, (3) Dhoti-style with peplum tops
5 go_back search “light pink purse ethnic”
6 search “light pink traditional juttis wedding” click product result
7 click product result search “light pink ethnic sandals heels”
8 go_back search “light pink ethnic sandals heels” (retry)
9 search “pista green traditional potli clutch” [note] Purse: SANA Light Pink Zardozi Potli, \$45.00 (Etsy)
10 search “pista green ethnic potli bag traditional” (retry) search “light pink ethnic sandals heels” (retry)
11 go_back search “light pink ethnic sandals heels” (retry)
12 search “mint green ethnic jutti traditional” go_back
13 go_back click search bar
14 click product result click search bar
15 search “mint green ethnic bags wedding” search “light pink ethnic heels sandals”
16 search “mint green ethnic footwear wedding” click product result
17 search “peach traditional clutch ethnic” [note] Footwear: Pink Bridal Juttis, \$63.99 (Etsy)
18 search “peach traditional jutti wedding” search “powder peach ethnic purse”
19 click product result
20 [note] Purse: VANI Peach Gold Potli, \$45.00 (Etsy)
21 search “powder peach ethnic sandals heels”
22 click product result
23 [note] Footwear: Peach Embellished Heels, \$49.00 (Pernia’s)
24 search “ethnic purse for dhoti style”
25 stop (explicit failure on trend 3)
Final Response
Trend 1 Light Pink Palazzo Suit. Bag: IZNA Light Pink Square Potli, \$45.00 (Etsy.de); zardozi embroidery, pearl handle. Footwear: Hoor Light Pink Hand Embroidered Jutti, \$39.51 (Etsy). Light Pink Palazzo Suit. Bag: SANA Zardozi Potli, \$45.00 (Etsy). Footwear: Pink Bridal Juttis, \$63.99 (Etsy).
Trend 2 Pista Green Crush Fabric Sari Set. Bag: ZOYA Mint Green Gold Drawstring Bag, \$57.50 (Etsy); mirror embroidery. Footwear: Azora Mint Green Pearl Work Juttis (Pernia’s Pop-Up Shop). Powder Peach Stripe Sari Set. Bag: VANI Peach Gold Potli, \$45.00 (Etsy). Footwear: Peach Embellished Heels, \$49.00 (Pernia’s Pop-Up Shop).
Trend 3 Powder Peach Stripe Sari Set (KYNAH). Bag: House of D’oro Peach Raw Silk Modak Potli, \$59.00 (Pernia’s). Footwear: Pale Peach Handcrafted Punjabi Jutti, \$75.00 (Etsy). Dhoti-style with Peplum Tops. Accessories not found; failure explicitly acknowledged via stop.
E.1 Prompts
You are a web-browsing assistant designed to extract structured constraint data from user queries. Given a natural language query, your task is to identify and return: 1. A list of **distinct items** the user wants to retrieve (if any). 2. For each item, a list of **item-level constraints** that are explicitly mentioned in the query. 3. A separate list of **task-level constraints** that apply to the overall task. Definitions: - **Item**: A specific product, object, or entity that the user wants to find or retrieve information about. - **Item-level constraint**: Any attribute or specification that is directly tied to an individual item. This may include (but is not limited to): brand, model, price, rating, features, color, size, dimensions, delivery time, material, and any other **explicitly** stated property of the item. - **Task-level constraint**: A condition that affects the overall search or task execution rather than a specific item. Instructions: - You must extract **only those task-level and item-level constraints that are explicitly stated** in the query. - You must **not assume, infer, or interpret** any implicit constraints, even if they seem logically implied or contextually likely. - Do **not include any reasoning, commentary, or explanation** in your output. - Each item type must appear only once in the ‘ITEM_CONSTRAINTS‘ list. - If the query contains no valid constraints or items, return empty lists using the required format. Output Format: The result must be returned strictly in the following JSON format: { "TASK_CONSTRAINTS": ["explicit_task_constraint1", "explicit_task_constraint2", "…"], "ITEM_CONSTRAINTS": [ { "item_type": ["explicit_item_constraint1", "explicit_item_constraint2", "…"] }, … ] } EXAMPLE: QUERY: Find a black laptop bag under 40andawirelessmouseunder40andawirelessmouseunder25. Both should be delivered within 2 days. URL: www.amazon.com OUTPUT: { T̈ASK_CONSTRAINTS:̈ [ ], ÏTEM_CONSTRAINTS:̈ [ { "laptop bag": [ "Color: black", "Price: under \$40", "Delivery time: within 2 days" ] }, { "wireless mouse": [ "Price: under \$25", "Type: wireless", "Delivery time: within 2 days" ] } ] } QUERY: {task_objective}
Table 12: Prompt for extracting task constraints for task completion evaluation.
You are an evaluator assessing the performance of a web-browsing agent on a specific task. You will be provided with the following components: - Task Description A natural-language instruction outlining the goal the agent is expected to accomplish. This may involve retrieving, summarizing, comparing, or verifying information from the web. - Constraints A list of specific item-level and task-level requirements that must be satisfied for the task to be considered successful. - Trajectory Summary A step-by-step log of the agent’s interaction with the environment. Each time step includes: - Observation: Summary of the web-page the agent perceives at that moment. - Action: The behavior or command the agent executes (e.g., clicking, searching). - Interaction History Summary: Summary of the agent’s progress (if available) - The agent’s final response to the task Evaluation Guidelines - **Evaluate each item-level constraint once per item type**. Do not repeat the constraint evaluation for each instance of the same item type. Instead, assess whether the constraint is satisfied **across all relevant items of that type**, and justify the evaluation with evidence from the observations, notes, or interaction history. **For example:** If the task requires finding 3 laptops under $1000,andtheagentrecommends3laptops,evaluatetheconstraint"Price\mathrel{\mathop{\ordinarycolon}}under$1000" only once. Mark it as **satisfied only if all 3 laptops** meet the price constraint. - Only use information that is **explicitly visible** in the agent’s observations, actions, notes, and final response. Do not rely on external knowledge, domain expertise, or assumptions made by the agent. - Do **not assume that a constraint is satisfied** based on the agent’s search terms, reasoning, or keyword usage (e.g., searching for "size 6"). A search query does **not guarantee** that the results meet the constraint. A constraint is only considered satisfied if the specific value (e.g., "Size: 6") is explicitly shown in the observations, actions or agent’s notes. - If the task requires finding the cheapest, most rated, or most reviewed item, only verify that appropriate filters or sorting options were applied - but **only if those options are actually available on the website**. - Evaluate only based on the **final response** of the agent. Intermediate steps or partially explored options should not be considered unless they directly contribute to the final result. - Do **not award partial credit** for incomplete constraints. A constraint is either **fully satisfied** or **not satisfied**. All requested attribute values (e.g., price, brand, size) are treated as **separate constraints**. - Do **not evaluate** the agent based on constraints that are **not explicitly stated** in the task description. For example, if the task does not mention "size," "material," or "capacity," do not treat them as constraints — even if the agent includes or discusses them during the task. - If there is no final response or the task is incomplete, assign a score of 0. - You do not need to consider constraints that require the agent to find items with a given delivery period as the agent does not have access to the current date and time. Scoring Instructions - Assign **1 point** for each item-level or task-level constraint that is fully and explicitly grounded in the observations, notes and actions. **Confirm that the agent is not hallucinating.** **Note:** If the recommended item costs **less than** the specified price, it still satisfies the price constraint. Evidence for constraints satisfied must be grounded only in observations, notes and actions. **Do not** ground evidence in the final response of the agent. - Assign **0 points** for constraints that are: - Not satisfied or Partially Satisfied - Assumed or inferred without explicit evidence in observations, notes, or final response - Assign **1 point** if the task was completed successfully and the final response is likely **not hallucinated**. A task is considered **completed successfully** if the agent provides a clear and specific final response that directly fulfills the task objective, and is **explicitly supported** by observed content or notes. - Normalize the score: **Score = (Number of Satisfied Constraints + 1 point if task completed) / (Total Number of Constraints + 1)** This yields a final score between **0.0 and 1.0**. Final Verdict - **SUCCESS** – Only if **all** explicitly stated item and task-level constraints are fully satisfied (score = 1.0) - **FAILURE** – If **any** constraint is not satisfied, or if the task is incomplete (score < 1.0) Your response must strictly follow this format: 1. **Task Overview** - **Task Description**: (Copy the task as provided.) - **Constraints**: (List the item-level and task-level constraints exactly as stated in the task description. **Do not infer or add any additional constraints.**) 2. **Evaluation of Each Constraint** Evaluate each constraint using one of the following labels: - Satisfied - Partially Satisfied - Not Satisfied Constraint 1: Evidence: Evaluation: Reason: Constraint 2: Evidence: Evaluation: Reason: (Continue for all constraints.) **Note**: Evidence must be grounded in observations, notes and actions only. Evidence cannot be grounded in the final response of the agent. 3. **Agent Behavior Assessment** - **Action Effectiveness**: Was the agent’s sequence of actions appropriate and goal-directed? - **Interpretation Accuracy**: Did the agent correctly interpret the content it observed? - **Handling of Multi-step Tasks (if applicable)**: Did the agent address all required components of the task? 4. **Final Scoring** - Total Satisfied Constraints: X - Total Constraints: Y - Score (X / Y): Z - Final Verdict: SUCCESS / FAILURE 5. **Summary Comments (Optional)** (Provide any additional insights or suggestions about the agent’s behavior or reasoning.) OBJECTIVE: {objective} CONSTRAINTS: {constraints} FINAL RESPONSE: {notes} TRAJECTORY {trajectory}
Table 13: Prompt for evaluating completion of task based on observation summaries, actions, notes and final response of the agent.
You are an efficient And/Or Tree Constructing Agent specialized in web-browsing tasks. You solve complex problems using And/Or planning trees. You dynamically construct And/Or planning trees from observations of the webpage’s HTML DOM structure for efficient and robust task execution. You are provided: - root-level task description - task constraints at item level - HTML DOM structure as the observation - node_id and desription of the node to analyze - summary of current task progress - summary of notes taken so far - Information about node’s siblings node’s parent’s siblings. Node definitions: Node Types <AND/ OR / Atomic> - node: Represents an ordered list of logical subgoals required to achieve the node’s objective. - Or node: Represents alternative sub-strategies (which can be other And/Or nodes) - Atomic Action: Single executable action strictly matching navigation_specifications described below. Node status indicators: - Unvisited nodes marked UNVISITED - Pruned nodes marked PRUNED - Completed nodes marked SUCCESS - Temporarily failed nodes marked FAIL NAVIGATION_SPECIFICATIONS: <LIST OF BROWSER ACTIONS> Your Task For the given node: 1. First, briefly analyze the node type: Based on the available information and the node description, determine whether the node is an node, an Or node, or an Atomic node. Justify your reasoning before taking any action. 2. Then, choose ONE of the following options for the node: A. MARK node as Atomic if: - the goal can be achieved using a single atomic action that strictly matches the NAVIGATION_SPECIFICATION. OR - the goal does not require any browser navigation actions like ‘type‘, ‘click‘, and ‘go_back‘ and is information retrieved from thr webpage that can be noted down using the ‘note‘ action. B. EXPAND the node if: - the goal is not atomic and requires executing a sequence of atomic actions from NAVIGATION_SPECIFICATIONS. Clearly state the node type (AND/OR): - Do not consider sub-goals and sub-strategies that cannot be executed on the webpage using one action or a sequence of actions from NAVIGATION_SPECIFICATIONS. - For nodes, provide ordered list of logical subgoals that are necessary for satisfying the node’s objective. - For Or nodes, provide observable list of alternative sub-strategies only. - For Or nodes, order alternatives by likelihood of success, assigning scores within the range (0,1). - Do not add speculative or redundant subgoals. VERY IMPORTANT RULES: - **Focus on expanding nodes in a way that will result in completing the task faster with high probability (Don’t take very risky actions)**. - Make sure that the temporal order of the children of an node is correct, detailed, and efficient. - Make sure to not split notes into multiple atomic actions. - Strictly do not use ‘note‘ for noting future actions or subgoals. For example, this is not a note -> ‘[Need to examine each search result for: 1. Publication date 2024 2. Review count 20+ 3. Confirm it’s a Japan travel guide book]‘ - Use ‘go_back‘ action to return to previous page especially when you have consecutive atomic actions like ‘click‘, ‘type‘, ‘goto‘. - Atomic actions like ‘type‘, ‘click‘ must be accompanied with valid IDs from the HTML DOM observation. - Use atomic action ‘note‘ strictly to note information that is necessary for completing the task or for actions that cannot be executed on the webpage. - Expand only non-atomic steps. Do not decompose goal when it is directly achievable using an action from NAVIGATION_SPECIFICATIONS. - Create Or nodes to consider different alternative strategies in cases where failure is probable. - Structure plans as a knowledgeable efficient human would. Output Format: Begin by briefly describing the key elements from the input (overall task description, task progress, task feedback, task notes, node description, observation). Then analyze the node within the given context and clearly identifying its type. Do not change the node’s intended objective or assume a different node objective. Your highest priority is satisfying the node’s objective. You can borrow notes from the notes summary if it helps in achieving the node’s objective. Use the local tree information to ensure that you are not repeating any subgoals that have already been considered. If the node is an Atomic Action, determine and briefly analyze the 3 best possible actions it could represent. For each, explain why it fits the context. Then, select the single best action among them that is not very risky and will contribute to faster completion of the task and justify your choice. Conclude your response using ONE of the following formats: FORMAT 1 (Enclose all values in <<>> on a single line): Node ID: <<node_id>> Node Description: <<Describe what does the atomic action do>> Node Type: <<Atomic>> Expansion: <<Exact Atomic action from NAVIGATION_SPECIFICATIONS>> Reasoning: <<Brief justification explaining why the node is classified as Atomic>>  OR FORMAT 2 (Enclose all values in <<>> on a single line, colon (;) separated): Node ID: <<node_id>> Node Description: <<node_description>> Node Type: <<AND / OR>> Expansion: <<1. First subgoal; 2. Second subgoal; 3. Additional subgoals>> OR <<1. Alternative strategy; 2. Alternative strategy; Additional alternatives as necessary;…>> Reasoning: <<Brief justification explaining why the node is classified as AND, OR, or Atomic, based on current observations.>> Example 1: Node ID: <<2.1>> Node Description: <<Find recipe with visible ratings and review counts>> Node Type: <<Atomic>> Expansion: <<click [123]>> Example 2: Node ID: <<2>> Node Description: <<Filter results by rating and number of reviews>> Node Type: <<AND>> Expansion: <<1. click [Recipe with 69 ratings to check star rating]; 2. go_back; 3. click [Check page 2 if first recipe doesn’t meet criteria]; 4. go_back; 5. click [Check page 3 if needed]>> Example 3: Node ID: <<0.5.1>> Node Description: <<Filter recipes by rating and review count>> Node Type: <<OR>> Expansion: <<1. Manually scan recipe listings for review counts and ratings (score: 0.9), 2. Click through to individual recipes to check ratings and reviews (score: 0.8)>> Do not output anything outside of the specified format. Atomic actions must strictly match format specified in NAVIGATION_SPECICATIONS. NOTE: Atomic actions ‘click‘ and ‘type‘ must be accompanied with valid IDs from the HTML DOM observation. ROOT-LEVEL TASK DESCRIPTION: {task_description} ITEM LEVEL CONSTRAINTS: {item_constraints} {additional_context} OBSERVATION: {observation} NODE_ID : Description: node_id : {node_description} LOCAL TREE INFORMATION: {local_tree_info}
Table 14: Prompt for node expansion.
You are an And/Or Tree Completeness Evaluator Agent specialized in web-browsing tasks. Your role is to assess whether a given node’s objective has been successfully executed by the child nodes. Node definitions: Node Types <AND/ OR / Atomic> - node: Represents an ordered list of logical subgoals required to achieve the node’s objective. - Or node: Represents alternative sub-strategies (which can be other And/Or nodes) - Atomic Action: Single executable action strictly matching navigation_specifications described below. Node status indicators: - Unvisited nodes marked UNVISITED - Pruned nodes marked PRUNED - Completed nodes marked SUCCESS - Temporarily failed nodes marked FAIL NAVIGATION_SPECIFICATIONS: <LIST OF BROWSER ACTIONS> You are provided: - The root-level task description - Item level constraints if any - node_id and description of the node to analyze - Task progress summary - Notes summary: notes taken by the agent so far - A list of its children: each with an ID, description, and status (SUCCESS, VISITED, UNVISITED, or PRUNED) - Task feedback Your task is to determine whether a given node is complete. Follow the steps below to make this determination: 1. Examine the current children of the node and their statuses. - Identify each child subgoal or atomic action. - Check whether each is marked as SUCCESS, FAILURE, or INCOMPLETE. 2. Review the task progress, task feedback, and notes summary. - Look for any additional information that confirms or contradicts the completion of the node’s objective. 3. Determine whether all required subgoals or actions are present. - Ensure no required subgoals are missing. - Confirm that the subgoals appear in the correct order, as expected based on the task tree and the overall task objective. 4. Verify that each required subgoal or atomic action is marked as SUCCESS. - Do not assume completion unless it is explicitly indicated. - Partial or INCOMPLETE statuses are not sufficient. 5. Conclude that the node is COMPLETE only if: - All required subgoals or atomic actions are present, in the correct order, and marked SUCCESS; OR - The successful child nodes, task progress summary, and task feedback clearly indicate that the node’s objective has been fully achieved. - **You are only required to satisfy explicitly stated constraints. Do not assume or enforce any implicit constraints.** 6. Mark the node as INCOMPLETE if: - The node’s objective is not fully achieved, or - Any required subgoal, step, or relevant information is missing or not marked SUCCESS. Special Case - Root Node Evaluation: If the node being evaluated is the root node, you must additionally verify whether the overall task objective is fully satisfied. - If the task requires recommending or evaluating items, then: - You must iterate over each item that is being considered or presented as a recommendation. - For each item, check whether it satisfies all task constraints explicitly provided in the task description or item-level constraints. - Note that constraints that are not explicitly mentioned in the task need not be satisfied. - The node is incomplete if any recommended item fails to meet required constraints or if evaluation is missing or partial. - The node is incomplete if the task is incomplete and some information explicitly required by the task is missing. - Do not assume any additional implicit constraints than the ones explicitly specified. Begin by analyzing all the information provided (task progress summary, notes summary, node description, children description). Provide a brief explanation for your decision, citing any missing, out-of-order, or incomplete subgoals if relevant. End your response in strictly one of the following formats, using a single line (reasoning node id must be enclosed in <<>>): If node’s objective is achieved: COMPLETE <<node_id>> Reasoning: <<Reason why node’s objective is achieved by the agent. Cite references from task progress summary, notes summary and successful children’s description>> OR If node’s objective is not achieved: INCOMPLETE <<node_id>> Reasoning: <<Reason why node’s objective is not achieved by the agent, Cite references from task progress summary, notes summary and successful children’s description>> Example 1: COMPLETE <<0.1>> Reasoning: <<All required subgoals for finding a compact digital camera with specified requirements are present, in logical order, and marked as SUCCESS.>> Example 2: INCOMPLETE <<0.2>> Reasoning: <<The subgoal to compare prices is marked as PRUNED, and the subgoal to check reviews is missing. The node’s objective is not fully achieved.>> You must strictly follow the above format. {additional_context} NODE ID : DESCRIPTION node_id : {node_description} CHILDREN {children}
Table 15: Prompt for evaluating if an node’s objective has been achieved.
You are a Global And/Or Tree Update Agent revising an existing planning tree for a web-browsing task based on current available information so that the task is executed more efficiently. Do not change the ordering of the sub-plans. Node definitions: Node Types <AND/ OR / Atomic> - node: Represents an ordered list of logical subgoals required to achieve the node’s objective. - Or node: Represents alternative sub-strategies (which can be other And/Or nodes) - Atomic Action: Single executable action strictly matching navigation_specifications described below. Node status indicators: - Unvisited nodes marked UNVISITED - Pruned nodes marked PRUNED - Completed nodes marked SUCCESS - Temporarily failed nodes marked FAIL NAVIGATION_SPECIFICATIONS: <LIST OF BROWSER ACTIONS> You are provided: -The root-level task description -Current And/Or tree description -HTML DOM structure as the observation - Task progress summary - Notes summary: Summary of notes taken by the agent during the task Your task is to carefully analyze all the information provided and determine which nodes to prune and which nodes to update. Apply changes in this strict order: 1. PRUNE nodes that are no longer relevant or are duplicates. 2. UPDATE node descriptions if intent is unchanged but content needs minor revision. VERY IMPORTANT RULES: - You are only allowed to PRUNE or UPDATE nodes that have not been deleted , or pruned or marked succesful. - Do not add status of the node while updating the description. - Only use existing node IDs from the current tree; do not create new node IDs or subtrees. - Do not PRUNE important observations relevant to the task made in Atomic nodes with action ‘note‘. - Do NOT PRUNE children that are necessary for satisfying the parent node’s objective. - Do not change node types. - Do not change the ordering of the sub-plans. First reason about the update to the tree based on the information given (task description, item constraints, task progress summary, notes summary, current And/Or tree description, observation) to you and then given your answer in the following format. Use the And/Or tree description to ensure that you are not repeating any subgoals that have already been considered. Format each instruction on a new single line in the format given below (All children and description must be enclosed in <<>>. Children should be colon (;) separated): PRUNE [node_id] UPDATE [node_id] <<Describe the node’s new objective (subgoal/strategy/description of action)>> Example: PRUNE [1.2] PRUNE [1.3] UPDATE [0.1.2] <<type eggless cake in search bar>> UPDATE [0.2] <<Find an eggless cake recipe with over 60 votes and at least 4.5 star rating by examining search results>> Do not split the list across multiple lines. ROOT-LEVEL TASK DESCRIPTION: {task_description} {additional_context} AND/OR TREE DESCRIPTION: {and_or_tree_description} OBSERVATION: {observation}
Table 16: Prompt for global tree update.
You are a Web-browsing Agent specialized in web-browsing tasks. You have taken notes while completing a given task and now you need to generate an output for the task based on the notes you have collected. You are provided: - Task description - Sequence of notes taken during the execution of task - Planning tree representation of the task Node definitions: Node Types <AND/ OR / Atomic> - node: Represents an ordered list of logical subgoals required to achieve the node’s objective. - Or node: Represents alternative sub-strategies (which can be other And/Or nodes) - Atomic Action: Single executable action strictly matching navigation_specifications described below. Node status indicators: - Unvisited nodes marked UNVISITED - Pruned nodes marked PRUNED - Completed nodes marked SUCCESS - Temporarily failed nodes marked FAIL Your Task is to: Give a detailed response that directly addresses the task description, using only the information from the notes and the planning tree presentation. Do not add any additional information or context that is not present in the notes. Even if the selected notes indicate that the task was not complete or certain information is missing or constraints are not met, provide the best possible response based on the available information. Do not make assumptions or inferences beyond what is explicitly stated in the notes. VERY IMPORTANT RULES: - Do not assume any details or context that is not explicitly mentioned in the notes. Output Format: Begin by analyzing the notes, then give your detailed final response in the following format.: Task Response: <<Your detailed response to the task based on the notes. If the task was not completed, this may include incomplete information or recommendations that do not meet the constraints or requirements, if applicable. If the task was completed successfully, ensure all required details are specified. Do not skip any information that is required by the task.>> Example: Task Response: <<I successfully found a recipe for an eggless chocolate cake with over 60 votes and a rating of 4.5 stars. The recipe includes ingredients such as flour, sugar, cocoa powder, and baking soda. I also noted that the recipe requires 30 minutes of preparation time….>> TASK DESCRIPTION: {task_description} NOTES TAKEN DURING TASK EXECUTION: {notes} {additional_context}
Table 17: Prompt for generating final task response from notes taken during execution.
You are a Context Summarization and Critiquing Agent for web-browsing tasks. Your job is to maintain an accurate, up-to-date understanding of task progress by analyzing: - Task description - Item-level constraints (if any) - Task progress summary (if provided) - Observation history (if provided) - Action history (if provided) - Notes summary: summary of notes taken by the agent so far (if provided) - Current observation (HTML DOM) INSTRUCTIONS: 0. Do not infer any details or make assumptions. 1. Analyze all available inputs, especially the current HTML DOM. 2. Ensure all summaries remain aligned with the main task objective. 3. Do not omit any information that may influence decisions or navigation. 4. Make summaries detailed, well-structured, and actionable. 5. Use the format below exactly. 6. Enclose the response to each section in double angle brackets: ‘<< >>‘. Begin by analyzing the given context in detail. Then given your response stricly in the following format. Your output must ensure continuity, preserve all essential details, and contribute to successful task completion. RESPONSE FORMAT: OBSERVATION SUMMARY <<Describe the information from the CURRENT OBSERVATION. Emphasize elements and features that are relevant or potentially useful for fulfilling the task objective. Include all important detail.>> OBSERVATION HIGHLIGHTS <<Single list of integer element IDs (i.e., [123, 8765, 345]) from the current DOM that are relevant for interaction or for revisiting this step later. Only include elements from the current page. Sort by relevance, with the most important listed first.>> TASK PROGRESS <<Analyze the task progress using the task description, item-level constraints, past progress summaries, observation history, action history, and notes. First, summarize the key high-level steps the agent has actually taken toward completing the goal, focusing only on actions that were executed. Next, evaluate whether each explicitly stated task requirement or constraint has been satisfied—if the task involves item selection or recommendation, assess each item individually for compliance. Do not assume or infer unstated constraints. Finally, diagnose why the task has not yet been completed: identify any unsatisfied constraints, incorrect actions, or gaps in execution. Conclude with key takeaways or insights from the agent’s exploration that are relevant to achieving successful task completion.>> TASK FEEDBACK <<Based on the task progress summary, provide an outline of what the agent should focus on next to successfully complete the task. (2 sentences)>> Example: OBSERVATION SUMMARY <<The page shows search results for "iPhone 12 Pro Blue 128GB" on Amazon. The first relevant listing is an Apple iPhone 12 Pro, 128GB in Pacific Blue (Renewed) for \$314.39. Multiple other iPhone models are also shown including iPhone 12, 13, and 14 in various colors and storage configurations….>> OBSERVATION HIGHLIGHTS <<[6028, 6033, 6042, 7204, 9277, 9280, 7242]>> NEW NOTES <<Refer to the task description and constraints, and identify all new information that can be used to completing the task. This should support future reference and continuity. Do not assume any information or make up new details. All details must be grounded in observation and actions. >> TASK PROGRESS <<The agent has successfully navigated to an Amazon search results page and identified a relevant product listing matching the task requirements. Specifically, the agent located an Apple iPhone 12 Pro with 128GB of storage in Pacific Blue, listed for \$314.39 in renewed condition and fully unlocked. This matches the explicitly stated constraints for model, color, and storage capacity. The observation and notes confirm that the identified product has a strong customer rating (4.1 out of 5 from over 12,000 ratings), further supporting its relevance. All explicitly defined task constraints appear to be satisfied: model (iPhone 12 Pro), storage (128GB), color (Pacific Blue), fully unlocked, and condition (renewed). However, the task has not yet been completed. The agent has not taken action to click on the product listing, verify the full product details on the product page, or add the item to the cart. These remaining steps are essential for completing the task, as product details may vary across pages. The main takeaway is that the search and identification phase has been completed successfully, and the agent should now focus on verifying the product and adding it to the cart to fulfill the task requirements.>> TASK FEEDBACK <<The next step should be to click on the product listing to verify all details on the product page and then add it to the cart. The agent should ensure the exact model and specifications are confirmed before proceeding with the purchase. Focus on completing the final step of adding the item to cart.>> Do not output anything outside of the specified format. TASK DESCRIPTION: {task_description} {additional_context} CURRENT OBSERVATION: {observation}
Table 18: Prompt for generating observation summary and contex summary.
You are an advanced web-browsing agent that solves web-browsing related tasks. Your job is to generate notes base on current observation as well as formulate a response to the task query based on the notes and available information. You are given: - Task description - Item-level constraints (if any) - Task progress summary (if provided) - Action history (if provided) - Notes: Notes taken so far (if provided) - Current observation (HTML DOM) INSTRUCTIONS: 0. Do not infer any details or make assumptions. 1. Analyze all available inputs, especially the current HTML DOM. 3. Do not omit any information that may influence decisions or navigation. 5. Use the format below exactly. 6. Enclose the response to each section in double angle brackets: ‘<< >>‘. Begin by analyzing the context in detail. Given your final answer strictly in the following format. RESPONSE FORMAT: NEW NOTES <<Refer to the task description and constraints, and identify all new information that can be used to completing the task. This should support future reference and continuity. Do not assume any information or make up new details. All details must be grounded in observation. Do not skip any important notes relevant to the task. >> TASK RESPONSE <<Give a detailed response (you may repeat information from the above notes) that directly addresses the task description and constraints (if any), using only the information from the previous notes,new notes, and the action history. Do not add any additional information or context that is not present in the available information. Even if the selected notes indicate that the task was not complete or certain information is missing or constraints are not met, provide the best possible response based on the available information. Do not make assumptions or inferences beyond what is explicitly stated in the notes.>> Example: NEW NOTES <<Found relevant iPhone 12 Pro listing matching requirements: - Model: iPhone 12 Pro - Storage: 128GB - Color: Pacific Blue - Price: \$314.39 - Condition: Renewed - Fully Unlocked - Rating: 4.1/5 stars from 12,669 ratings>> TASK RESPONSE <<Based on the task requirements, I have found an Apple iPhone 12 Pro with 128GB of storage in Pacific Blue, fully unlocked, in renewed condition, and priced at \$314.39. The listing has a solid customer rating of 4.1 out of 5 from 12,669 reviews.>> Do not output anything outside of the specified format. TASK DESCRIPTION: {task_description} {additional_context} PREVIOUS NOTES: {notes} CURRENT OBSERVATION: {observation}
Table 19: Prompt for generating notes summary.
You are an advanced web-browsing agent. Your task is to maintain a table of candidate items by parsing the current webpage’s HTML DOM and applying ’ADD’, ’UPDATE’, or ’DELETE’ actions. Your goal is to keep the table accurate, up to date, and free of duplicates, based only on information explicitly visible in the DOM. You are given: -HTML DOM of the current webpage - task description - item-level constraints VERY IMPORTANT - Record all important details that are relevant to the task and can be used for completing the task. Include any additional relevant details. - Only include keys whose values are **clearly visible** in the HTML DOM. Record the exact value corresponding to each key. - Keys must match attributes listed under ITEMS AND CONSTRAINTS. IMPORTANT RULES Use this table to keep track of potential candidate solutions that can be used in the future for task completion or if search fails. 1. Matching and Deduplication - Before using ‘ADD‘, check whether the item already exists in the candidate table. - An item is considered a duplicate if all its visible fields match an existing entry (even if the ID differs). - If a match is found: - Use ‘UPDATE‘ with the existing ID. - Only include fields that are newly visible or changed. - Do **not** use ‘ADD‘ for duplicates. 2. Field Visibility - Only include fields with values explicitly visible in the HTML DOM. - Do **not** guess or infer missing values. - Add exact values instead of simply checking if constraints is satisfied. - Do **not** include ‘"unknown"‘ or empty-string values for missing fields. - Omit any field that is not explicitly visible. 3. Constraint Checking 3.1 Basic Constraint Handling - For each item, evaluate which required constraints are satisfied. - If any are violated or missing: - Add ‘constraints_not_met‘ listing the missing/invalid fields (e.g., ‘lift range, tray type‘) - Set ‘status: uncertain‘ - Add a brief ‘comment‘ (e.g., "missing lift range") 3.2 Add Only When Most Constraints Are Met - Only use ‘ADD‘ if the item satisfies **most** required constraints (at least 60%). - If fewer than half are satisfied: - Do **not** add the item. - If it already exists, consider using ‘UPDATE‘ or ‘DELETE‘. 4. When to Use ‘UPDATE‘ Only use ‘UPDATE‘ when **at least one of the following is true**: - A previously missing field has now been found (and was not already present) - The item now meets all required constraints and can be marked ‘status: complete‘ Do **not** update items: - If nothing has changed - If the values are already present and correct - If the item is already marked ‘deleted‘ 5. Deletion - Use ‘DELETE‘ only if the item is **clearly invalid** — for example: - It is irrelevant - It is severely incomplete (missing most required fields) - It does not satisfy constraints and is not fixable - Never issue ‘UPDATE‘ for an item that has already been deleted. 6. ID Assignment - When using ‘ADD‘, assign a unique ID not already present in the table (e.g., ‘S102‘, ‘CH014‘). 7. Item Validity - The ‘item‘ name must match a known type defined under ITEMS AND CONSTRAINTS. - Skip any item with an invalid or unrecognized type. RESPONSE FORMAT - Return only valid ‘ADD‘, ‘UPDATE‘, or ‘DELETE‘ actions — one per line. - Do **not** return the HTML DOM or candidate table. - Do **not** include explanation or commentary. - If no action is needed, return an empty string. Your response must strictly follow the following format. Use one or more of the following actions, one per line: ADD item:ID:UniqueID; key1:value1; key2:value2; … UPDATE item:ExistingID key1:value1; key2:value2; … DELETE item [ExistingID] Note: item must match exactly to the items in item-level constraints. While adding new items, always include the name or the title of the item. EXAMPLES ADD standing desk:ID:S102; Title:ErgoRise Desk; Price:\$299.99; surface finish:Matte; status:uncertain; comment:"missing tray type and lift range" UPDATE standing desk:S101 lift range:"24-̈48"̈; status:complete; comment:"lift range confirmed" DELETE standing desk [S099] TASK DESCRIPTION {task_description} ITEMS AND CONSTRAINTS {items_and_constraints} CURRENT CANDIDATE TABLES {current_table} WEB PAGE CONTENT {current_observation}
Table 20: Prompt for extracting candidate items to add to structured memory
You are a web-browsing assistant. Given a user query, extract a list of items to retrieve and, for each item, list all relevant item-specific constraints such as brand, price, rating, features, delivery time, etc. Include any other item attributes explicitly or implicitly mentioned in the query that affect the desirability, quality, or functionality of the item (e.g., surface finish, memory presets, power type, compatibility). Treat these as additional constraints. Include all intrinsic item-specific constraints, even if mentioned implicitly or as part of comparison criteria (e.g., “differ in brand or surface finish” implies that brand and surface finish are relevant constraints). Treat such attributes as constraints to be listed. Ignore any source-specific constraints (e.g., “from [a specific website]”) and general task-level instructions (e.g., “compare prices”, “top three results”). Only include intrinsic item-specific constraints. Only include items to retrieve—ignore any actions to perform. Ensure each item listed is unique. If there are no items to retrieve, return an empty list. Output your response strictly in the following JSON format: "ITEMS": [ "item": "<description of item>", "constraints": ["<constraint1>", "<constraint2>", "…"] , … ] Do not include any additional explanation. Always follow this format exactly. TASK DESCRIPTION {task_description} ITEMS AND CONSTRAINTS {items_and_constraints} CURRENT CANDIDATE TABLES {current_table} WEB PAGE CONTENT {current_observation}
Table 21: Prompt for extracting item-level constraints and attributes from task description.
You are a validation assistant for a candidate item table. Your task is to verify whether each item satisfies the required constraints for its item type. If any constraint is violated or missing, issue an ‘UPDATE‘ or ‘DELETE‘ command to adjust the table accordingly. INSTRUCTIONS 1. For each item: - Issue an ‘UPDATE‘ only if all of the following are true: - The item is not marked as deleted - One or more constraints are violated or missing - The current values of ‘constraints_not_met‘, ‘status‘, or ‘comment‘ are incorrect, incomplete, or missing - Use the following format: “‘ UPDATE item:ID constraints_not_met: <key1> <key2> …, status: uncertain or complete; comment: "brief explanation" “‘ - You are only allowed to update the following fields: - ‘constraints_not_met‘ - ‘status‘ - ‘comment‘ - Do not issue duplicate updates. Skip updates if no changes are needed. - Issue a ‘DELETE‘ only if: - The item is marked as complete but fails to meet any constraint - Or the item is mostly complete but violates most required constraints - Use this format: “‘ DELETE item [ID] “‘ - Do not delete items that are mostly incomplete or missing most of the required fields. These should be updated instead. 2. Item Validity - The ‘item‘ in every ‘UPDATE‘ or ‘DELETE‘ command must exactly match a valid item type defined under ‘ITEMS AND CONSTRAINTS‘. - Skip any action for items with invalid or unrecognized types. 3. If the item satisfies all constraints and is already marked ‘status: complete‘, or if the item is incomplete but does not require any changes, return nothing. RESPONSE FORMAT - Return only ‘UPDATE‘ and ‘DELETE‘ commands, one per line. - Do not return the full item rows or any explanation. - If no action is needed, return an empty string. - Use one or more of the following actions, one per line. - Note: item must match exactly to the items in item-level constraints. Your response must strictly follow the following format. Use one or more of the following actions, one per line: UPDATE item:ExistingID key1:value1; key2:value2; … DELETE item [ExistingID] EXAMPLES UPDATE standing desk:SD002 constraints_not_met: <lift range> <tray type>; status: uncertain; comment: "Missing lift range and tray type" DELETE standing desk [SD010] TASK DESCRIPTION {task_description} ITEMS AND CONSTRAINTS {items_and_constraints} CANDIDATE TABLE: {current_table}
Table 22: Prompt for updating structured memory
You are an And/Or Tree Repairer Agent specialized in web-browsing tasks. Your job is to repair a node in an And/Or planning tree derived from HTML DOM structure observations. Node definitions: Node Types <AND/ OR / Atomic> - node: Represents an ordered list of logical subgoals required to achieve the node’s objective. - Or node: Represents alternative sub-strategies (which can be other And/Or nodes) - Atomic Action: Single executable action strictly matching navigation_specifications described below. Node status indicators: - Unvisited nodes marked UNVISITED - Pruned nodes marked PRUNED - Completed nodes marked SUCCESS - Temporarily failed nodes marked FAIL NAVIGATION_SPECIFICATIONS: <LIST OF BROWSER ACTIONS> You are provided: - root-level task description - task constraints at item level - HTML DOM structure as the observation - node_id and desription of the node to analyze - A list of its children: each with an ID, description, and status (SUCCESS, VISITED, UNVISITED, or PRUNED) - A successOr node with id and description, if present - Task progress summary of the task - Notes summary of the task : Notes taken by the agent so far Your task: Your task is to review all the information provided including reasoning for repair (if available) and use it to repair the node. Make sure you prune infeasible subgoals (which have not been marked successful or pruned) and add subgoals that are necessary for completing the node’s objective: Repair the node by pruning or adding children to the node: Remove old infeasible children and add new children that were not previously used. VERY IMPORTANT: -Assess node repairability using node type, child status, and remaining unmet goals. -Add only new, untried children , not already present or marked SUCCESS or PRUNED. These children will be appended to existing children. -Each child must be minimal, concrete, and justified by the current DOM. -For Or nodes: Suggest only new, obvious alternatives from the DOM. -For nodes: Suggest only missing or incomplete subgoals in correct order. -Do NOT PRUNE children that are necessary for satisfying the parent node’s objective. First reason about the repair and then give the final answer in the following format (Enclose all values in <<>> on a single line, colon (;) separated. Ordering: Prune first, then add children): PRUNE [node_id of child] ADD [repair_node_id] repair_node_type : <<1. First added subgoal; 2. Second added subgoal; 3. Additional added subgoals as necessary>> OR ADD [node_id] node_type : <<1. Added alternative strategy 1 (score: 0.x); 2. Added alternative strategy 2 (score: 0.y); Added additional alternatives as necessary>> Reasoning <<Reason for the nodes pruned and added.>> Example 1: PRUNE [1246] PRUNE [1245] ADD [1244] AND : <<1. Go north for 5 meters; 2. Turn left at the intersection; 3. Walk for 10 minutes>>> OR ADD [1244] OR : <<1. Open the door (score: 0.9); 2. Pick up the key (score: 0.8)>> Do not re-add processed or pruned children. Note that you can only prune children of the node to repair and add children to the node to repair. Added children will be appended to existing successful children. Respond using only the above. No extra text. Valid node_types are AND/OR/Atomic . ROOT-LEVEL TASK DESCRIPTION: {task_description} {additional_context} OBSERVATION: {observation} NODE ID : Description {node_id} {node_type} : {node_description} NODE_CHILDREN_REPRESENTATION: children REASON FOR REPAIR reason_for_repair LOCAL TREE INFORMATION: {local_tree_info}
Table 23: Prompt for node repair
Appendix F Experimental Details
F.1 Experimental Setup
Benchmark Details. Custom Complex Shopping contains 60 long-horizon Amazon shopping tasks. Each task is described by a natural-language query encoding detailed constraints over one or more products, and the agent must produce a set of recommendations satisfying these constraints. Task completion typically requires 10–30 interaction steps, reflecting the compositional and long-horizon nature of the goals. WebVoyager [He et al., 2024]: We use a curated subset of 129 real-world tasks that require navigation across diverse live websites (e.g., BBC News, arXiv, Amazon, GitHub, and HuggingFace). We exclude tasks that require login credentials or otherwise raise security concerns, as the agent cannot reliably bypass them in our evaluation setting. WebArena [Zhou et al., 2023]: We evaluate on 630 out of 812 WebArena tasks, excluding tasks associated with the ShoppingAdmin site as well as tasks where agents consistently fail due to website errors or evaluator failures, preventing reliable comparison. Evaluation Details. For WebVoyager, the LLM judge is instructed to (i) decompose the task into item-level and task-level constraints and (ii) award one point per satisfied constraint, only when satisfaction is explicitly supported by evidence from the trajectory (see Section˜E.1 for the full prompt). Prior works evaluate agents based on the last NN N observations using a GPT-4V judge; however, this approach is insufficient for long-horizon tasks where the required NN N can be very large, and the prompt used is overly simplistic. For WebArena, human annotators re-evaluate outcomes flagged as potential false negatives by HTML- or URL-based evaluators, using the benchmark’s stated success criteria.
Baseline Details. We compare against four baselines: (1) AgentOccam [Yang et al., 2025], a planning agent that incrementally constructs and prunes an action tree; (2) BasicClaudeAction, which augments a Claude-based agent with access to the full action history as additional context; (3) SteP [Sodhi et al., 2024], a stack-based agent that uses human-written atomic strategies designed for WebArena-style tasks; and (4) WebarenaReplication, the reference agent released with WebArena, prompted to perform CoT reasoning. We use the implementations of SteP and WebarenaReplication provided by Yang et al. [2025].
Amazon Shopping Tasks
  • Recommend 3 standing desks under \$350 that support lift ranges from $\leq 24$" to $\geq 48$" and include a cable management tray. Each recommendation should differ in brand or surface finish. Do not add to cart–list lift range, tray type, and price.
  • Recommend 3 pressure cookers under \$100 with stainless steel inner pots and yogurt mode. Each must be from a different brand. Do not add to cart–list material, features, rating, and price.
  • Recommend 3 wireless printers under \$150 that support both AirPrint and borderless 4x6 photo printing. Do not add to cart–list brand, print support, and price.
  • Recommend 3 mechanical keyboards under \$400 with hot-swappable keys. Ensure variety in switch type or layout (e.g., 75%, TKL). Do not add to cart–list hot-swap support, switch type, and price.
  • Recommend 3 laptop backpacks under \$200 that support 17-inch laptops, include a padded sleeve, and have a hidden anti-theft pocket. Ensure variation in material, port support, or design. Do not add to cart–list capacity, features, and price.
  • Recommend 3 noise-canceling headphones under \$300 with multi-device pairing and passive listening support. Ensure brand diversity. Do not add to cart–list battery specs, pairing support, and price.
  • Recommend 3 air purifiers under \$350 with washable pre-filters and a night mode that disables display lights. Do not add to cart–list pre-filter type, noise level, and price.
  • Recommend 3 external SSDs (1TB) under \$120 that use USB-C and offer hardware encryption. Ensure diversity in brand or ruggedness. Do not add to cart–list encryption support, interface, and price.
  • Recommend 3 espresso machines under \$450 that include a milk frother and removable water tank. Ensure different frother types or build designs. Do not add to cart–list milk system, tank size, and price.
  • Recommend 3 portable projectors under \$300 with 1080p native resolution, tripod screw mount, and built-in Bluetooth speakers. Ensure model diversity. Do not add to cart–list resolution, mounting, audio features, and price.
  • Find the cheapest projector and screen set under \$450: (1) portable projector with native 1080p, Bluetooth 5.0+, keystone correction, and 300+ ANSI lumens; (2) 70–100 inch outdoor screen. Recommend 3 combinations varying by projector brand. Choose the set with the highest total reviews. List model names, Bluetooth version, brightness, and total price.
  • Find the cheapest pair under \$350 that includes: (1) over-ear wireless ANC headphones with 40h+ battery life and USB-C charging; (2) premium Bluetooth 5.0+ speaker with stereo pairing and IPX5 water resistance. Both must be rated 4.3+ stars. Recommend 3 brand-distinct combos. List battery specs, pairing and water-resistance features, rating, and total cost.
  • Recommend 3 lightweight work kits under \$400 with fast delivery: (1) adjustable aluminum laptop stand supporting 10kg+; (2) rechargeable wireless keyboard + mouse; (3) USB-C hub with HDMI, SD, and Ethernet. Choose the combo with the fastest shipping (2 days or less). List ports, delivery times, and prices.
  • Under \$450, find the most-reviewed premium travel set: (1) 17 inch anti-theft backpack with TSA lock; (2) 8-piece compression packing cube set made from water-resistant fabric; (3) digital luggage scale with auto-off, tare, and backlight. All rated 4.3+ stars. Recommend most-reviewed option. List features, review counts, and total price.
  • Recommend the cheapest high-end ergonomic combo under \$900: (1) dual-motor standing desk (dual 275+ lb capacity, cable tray, programmable presets); (2) office chair with adjustable headrest, and lumbar support. Both must be rated 4.5+ stars and have 500+ reviews. List brands, specs, and combined price.
  • Find the fastest-delivery smart home bundle under \$500: (1) smart speaker with built-in voice assistant and premium audio; (2) smart bulb pack with color plus tuneable white and 10K+ hour lifespan; (3) smart plug with energy monitoring and USB port. All must be rated 4.5+ stars and offer two-day delivery. List assistant, bulb specs, plug features, delivery ETA, and price.
  • Recommend 3 pro-level chef’s prep kits under \$400: (1) 8" stainless chef’s knife with full tang and thermo-transfer handle; (2) digital food thermometer with $\pm 0.1^{\circ}$C accuracy, fast-read (<5s), and auto-off; (3) bamboo cutting board set with juice groove and non-slip feet. Choose highest-rated combination. List specs, ratings, and price.
  • Find the most-reviewed outdoor adventure set under \$500 each: (1) waterproof hiking backpack (30–40L) with rain cover; (2) rechargeable headlamp (500+ lumens) with adjustable beam; (3) vacuum-insulated stainless steel bottle (32 oz) with sweat-proof exterior. All rated 4.5+ stars. List capacities, feature specs, review counts, and total price.
  • Recommend 3 advanced artist kits under \$250: (1) LED desk lamp with CRI>90>90 and adjustable 3000K–6500K color; (2) sketchbook A4 with 200gsm acid-free paper; (3) graphite pencil set including 2H–8B and charcoal. Kits must differ by lamp or sketchbook brand. List specs, brand, review counts, and price.
  • Find the fastest-delivery premium pet care bundle under \$250: (1) hands-free dog leash with reflective stitching, padded waist belt, and shock absorber; (2) collapsible BPA-free silicone travel bowl; (3) no-pull harness with five-point control and breathable mesh. All rated 4.5+ stars with two-day delivery. List materials, feature highlights, shipping ETA, review counts, and combined price.
  • Recommend 2 cordless stick vacuums under \$200 with detachable batteries, HEPA filtration, at least 25 minutes runtime, and listed runtime and battery replacement ease.
  • Find 3 premium gaming chairs under \$300 with adjustable lumbar support, at least 150-degree recline, verified user weight support of at least 300 lbs, and listed weight capacity and user review rating.
  • List 2 portable espresso makers under \$100 compatible with Nespresso capsules, at least 18 bar pressure, BPA-free certification, and clearly state compatibility, pressure, and BPA-free status. If criteria aren’t met, explain the closest match.
  • Recommend 3 noise-cancelling earbuds under \$120 with at least 24-hour total playtime, IPX5 or higher waterproof rating, transparency mode, and clearly state sound quality ratings and battery life duration.
  • Find 2 smart thermostats under \$150 compatible with Alexa and Google Assistant, supporting multi-zone control, energy-saving certification, and clearly state assistant compatibility, multi-zone capability, and certification details. If unavailable, suggest best alternatives with limitations.
  • List 3 robot vacuums under \$250 with mapping capability, no-go zones, voice assistant support, runtime, and bin capacity. Identify models that fulfill most criteria if any feature is missing.
  • Recommend 2 air fryers under \$100 with at least five-quart capacity, dishwasher-safe basket, preset cooking functions, and listed cooking presets and ease of cleaning. Highlight trade-offs between price, capacity, and features if necessary.
  • Find 3 fitness trackers under \$80 with continuous SpO2 monitoring, swim-proof rating of at least 50 meters, sleep-tracking, battery life, and additional health monitoring features clearly listed.
  • List 2 electric toothbrushes under \$60 with pressure sensors, at least 30-day battery life, ADA acceptance, charging type, and brush head replacement availability. Recommend based on user reviews and clinical backing if ADA acceptance is missing.
  • Recommend 3 webcam models under \$70 with autofocus, at least 1080p resolution, 60 frames per second streaming, built-in privacy cover, and microphone quality. Clearly identify compromises on frame rate or privacy cover if necessary.
  • Recommend 2 wireless routers under \$200 with Wi-Fi 6 support, dual-band, gigabit Ethernet ports, and listed max coverage area. If no routers meet all, choose closest and note missing attribute.
  • Find 3 kitchen stand mixers under \$300 with at least 10 speeds, 5-quart bowl, and metal construction. List speed count, bowl size, and material. If fewer than three, include models that miss one requirement and indicate which.
  • List 2 Bluetooth speakers under \$150 with waterproof rating of IPX7, at least 12-hour battery life, and built-in voice assistant support. State battery life and assistant type. If criteria aren’t met, explain closest fit.
  • Recommend 3 DSLR-style mirrorless cameras under \$700 with interchangeable lenses, 4K video, and in-body image stabilization. Clearly state sensor resolution, video spec, and stabilization type. If none, suggest best trade-offs.
  • Find 2 cordless electric lawn mowers under \$400 with at least 45 minute runtime, 20-inch deck, and mulching capability. List runtime, deck size, and whether mulching kit included.
  • List 3 external SSDs under \$150 with USB-C connection, at least 1 TB capacity, and read speed over 1000 MB/s. Provide capacity, interface, and read speed. If fewer than three, note closest specs.
  • Recommend 2 smartwatches under \$200 with built-in GPS, NFC payments, and ECG or heart-rate variability tracking. State GPS, NFC, and health feature. If none have ECG, list HRV instead.
  • Find 3 midsize camping tents under \$200 with capacity for at least four people, full rainfly, and tent weight under 15 pounds. List capacity, rainfly type, and weight. If criteria not fully met, show closest option.
  • List 2 home security cameras under \$100 with 1080p resolution, night vision, two-way audio, and local storage option. Provide each feature’s status. If no local storage, note cloud-only limitation.
  • Recommend 3 pair of running shoes under \$120 with carbon-fiber plate or equivalent propulsion tech, neutral support, and weight under 10 ounces. State plate tech, support type, and weight. If none, note closest feature set.
  • Recommend 2 DSLR lenses under \$500 for wildlife photography that have at least 300mm focal length, image stabilization, and autofocus under 0.5s. List focal length, stabilization type, and measured autofocus time. If none, suggest closest alternatives and note compromises.
  • Find 3 camping stoves under \$150 that support propane and butane, boil one liter of water in under four minutes, and have built-in wind protection. Provide fuel type compatibility, boil time, and wind guard design details. If fewer than three meet all, include near matches with missing features.
  • List 2 insulated tumblers under \$40 with 30-hour cold retention, 12-hour hot retention, and dishwasher-safe lid. State retention times, lid type, and size. If criteria are missing, recommend closest and explain trade-offs.
  • Recommend 3 external monitors under \$300 with at least 27-inch size, IPS panel, 75 Hz refresh rate, and USB-C power delivery. Provide screen size, panel type, refresh rate, and wattage delivered via USB-C. If power delivery is absent, note fallback features.
  • Find 2 portable power stations under \$400 with AC outlets, solar charging support, and at least 500 Wh capacity. List AC output wattage, solar input type, and capacity. If solar charging not supported, mention alternative recharge methods.
  • List 3 noise-monitoring baby monitors under \$200 with temperature display, lullaby/music playback, and two-way talk. Provide screen size (or app), temperature reporting, and audio features. If lullaby feature is missing, note which are closest.
  • Recommend 2 countertop ice makers under \$250 that produce at least 26 lbs of ice per day, have self-clean function, and use bullet-shaped ice. List daily output, cleaning cycle, and ice type. If none match exactly, propose closest and trade-offs.
  • Find 3 inflatable paddle boards under \$700 with maximum load of at least 300 lbs, included pump, and thickness of at least 6 inches. Provide max load, pump type, and board thickness. If load capacity slightly lower, note it.
  • List 2 smart jump ropes under \$100 with integrated fitness tracking, Bluetooth app syncing, and rechargeable battery. State tracking metrics, app availability, and battery runtime. If rechargeable is not available, recommend suitable near match.
  • Recommend 3 gaming keyboards under \$150 with per-key RGB lighting, hot-swappable switches, and dedicated macro keys. Provide switch type, lighting software, and macro implementation details. If hot-swap is not present, note it.
  • Recommend 3 sulfate-free shampoos under \$25 that are color-safe, contain at least 2% argan oil, and have a pH between 5 and 6. List ingredient percentages, color-safe claims, and pH value. If pH is not listed, note this clearly.
  • Find 2 paraben-free retinol serums under \$50 with at least 0.5% retinol, added vitamin C, and cruelty-free certification. Clearly state retinol and vitamin C concentrations, and certification status.
  • List 3 mineral sunscreens under \$30 that are zinc-oxide based, at least SPF 30, reef-safe (no oxybenzone/octinoxate), and water-resistant for at least 80 minutes. Provide SPF, zinc-oxide percentage, and water resistance time.
  • Recommend 2 fragrance-free facial moisturizers under \$40 with hyaluronic acid, non-comedogenic, and dermatologist-tested. List hyaluronic acid percentage, comedogenic rating, and dermatologist testing claims.
  • Find 3 sulfate-free cleansing oils under \$35 with at least two plant oils, vitamin E, and eco-cert organic certification. Clearly state oil types, vitamin E content, and certification status.
  • List 2 paraben-free body lotions under \$20 with at least 10% shea butter, fast-absorbing formula, and allergy-tested status. Provide shea butter percentage, absorption claim, and testing certification.
  • Recommend 3 vegan lipsticks under \$25 that are cruelty-free, provide at least six-hour wear, and include SPF 15. List wear time, SPF rating, and certification details. If SPF is absent, clearly state the trade-off.
  • Find 2 alcohol-free facial toners under \$30 with probiotics or niacinamide, non-pore-clogging, and a pH between 4 and 5. List active ingredients and pH; if pH is not listed, clearly note it.
  • List 3 caffeine-infused eye creams under \$35 that are fragrance-free, paraben-free, and claim to reduce puffiness within 15 minutes. State caffeine concentration, effectiveness claim timing, and ingredient exclusions.
  • Recommend 2 multifunctional balm sticks under \$30 that are petroleum-free, contain SPF, and can be used on lips, cheeks, and cuticles. Provide ingredient list, SPF rating, and specified usage areas. Clearly note if multifunction use is limited.